### v2

# Test-time Correction with Human Feedback: An Online 3D Detection System via Visual Prompting

Abstract
This paper introduces Test-time Correction (TTC) system, a novel online 3D detection system designated for online correction of test-time errors via human feedback, to guarantee the safety of deployed autonomous driving systems. Unlike well-studied offline 3D detectors frozen at inference, TTC explores the capability of instant online error rectification. By leveraging user feedback with interactive prompts at a frame, e.g., a simple click or draw of boxes, TTC could immediately update the corresponding detection results for future streaming inputs, even though the model is deployed with fixed parameters. This enables autonomous driving systems to adapt to new scenarios immediately and decrease deployment risks reliably without additional expensive training. To achieve such TTC system, we equip existing 3D detectors with Online Adapter (OA) module, a prompt-driven query generator for online correction. At the core of OA module are visual prompts, images of missed object-of-interest for guiding the corresponding detection and subsequent tracking. Those visual prompts, belonging to missed objects through online inference, are maintained by the visual prompt buffer for continuous error correction in subsequent frames. By doing so, TTC consistently detects online missed objects and immediately lowers driving risks. It achieves reliable, versatile, and adaptive driving autonomy. Extensive experiments demonstrate significant gain on instant error rectification over pre-trained 3D detectors, even in challenging scenarios with limited labels, zero-shot detection, and adverse conditions. We hope this work would inspire the community to investigate online rectification systems for autonomous driving post-deployment. Code would be publicly shared.
Figure 1: TTC system could instantly respond to missed objects via human feedback, detecting and tracking these targets in subsequent frames consistently. As depicted on the left, human feedback might come from any views of target objects, including missed objects in the current and/or previous frames, as well as diverse viewpoints across scenes, styles, sources, and poses. Such a system improves offline-trained 3D detectors by rectifying online driving behavior immediately, reducing safety risks via test-time correction. TTC enables a reliable and adaptive online autonomous driving system.
Equal contribution.
1 Introduction
Figure 2: Comparison of Error Correction between the conventional offline loop (left) and the new proposed online TTC System (right). Offline error correction pipeline improves model capability during the development stage, which typically requires expensive workloads and computational overhead over days or weeks for model updates. While TTC system additionally enables deployed 3D detectors with on-the-fly error rectification ability.
Visual-based 3D object detection, which localizes and classifies 3D objects from visual imagery, plays a crucial role in autonomous driving systems. Visual autonomous driving frameworks [20, 19, 3, 9] rely heavily on accurate 3D detection outcomes to predict future driving behaviors and plan the trajectory of the ego vehicle. Existing 3D object detectors [27, 62, 35, 46, 21] typically follow an offline training and deployment pipeline. Once the model is trained and deployed on self-driving cars, it is expensive to update new behaviors, e.g., another turn of offline re-training or fine-tuning. That is, when the system fails to perceive important objects or fails in novel scenarios due to a domain shift, these offline solutions cannot update themselves online to rectify mistakes immediately and detect missed objects. Such a caveat poses significant safety risks to reliable driving systems, e.g., dangerous driving behaviors such as improper lane changing, turning, or even collisions.
To guarantee safety, we argue that 3D detectors deployed on autonomous driving systems can rectify missed detection on the fly during test time. As depicted in Figure 2 (left), for error rectification, existing 3D detectors rely on offline pipelines, encompassing a full-suite procedure of data collection, annotation, training, and deployment. This requires significant human workloads and resources for labeling and re-training, days or even weeks to fulfill. Other than the offline pipeline to improve models, we desire deployed 3D detectors also capable of test-time correction since such delays in offline updating are unacceptable when facing risks on the road, where safety is of the utmost priority.
In this work, we explore a new 3D detection system capable of Test-Time error Correction based on human feedback online, namely TTC, akin to how human drivers respond, as shown in Figure 2 (right). The system is designed to address missed detection issues in autonomous driving scenarios, enabling existing 3D detectors to instantly correct errors through online human input, thereby preventing the risks of erroneous decisions caused by missed detections. Inspired by the principles of In-context Learning Customization [58, 43] in large language models (LLMs) [40, 52], we achieve the TTC system by leveraging images of missed objects collected from human feedback as context prompts. These prompts assist deployed 3D detectors in identifying and localizing previously unrecognized objects in later streaming input frames, without extra parameter updates during testing.
The proposed TTC includes two components: Online Adapter (OA) that enables 3D detectors with visual promotable ability, and a visual prompt buffer that records missing objects. The core design is “visual prompts”, the visual object representation derived from human feedback. Existing promptable 3D detection methods typically utilize text, boxes, or clicks as prompts [4, 7, 66, 15, 2]. However, text prompts can be ambiguous and may not describe the target objects effectively. Meanwhile, box and point prompts struggle to handle streaming data. These limitations indicate that such prompts are inadequate for real-time autonomous driving tasks. Visual prompts cover arbitrary imagery views of target objects, i.e., views in different zones, styles, and timestamps (Figure 3), indicating the identity of target objects. With visual prompts, OA module generates corresponding queries, locates corresponding objects within streaming inputs, and facilitates 3D detectors to output 3D boxes.
To enable consecutive error rectification for video streaming, we design a dynamic visual prompt buffer to maintain visual prompts of all past unrecognized objects. In each iteration, we use all visual prompts in this buffer as inputs of the TTC system. This enables the continuous detection and tracking of all previously missed objects, redressing online errors for streaming input effectively. Further, to refrain the buffer from undesirable expansion, we introduce a “dequeue” operation to ensure its bounded size, allowing for consecutive rectification without excessive overhead.
Figure 3: Visual prompts could be arbitrary views of objects, across zones, styles, timestamps, etc.
We conduct experiments on the nuScenes dataset [1]. Given the novel setting of our TTC system, the assessment focuses on its abilities for instant online error correction. To this end, we design extensive experiments to verify the key aspects of the overall system: the effectiveness of TTC system in rectifying errors over the online video stream; and the effectiveness of TTC when encountering challenging, extreme scenarios with large amounts of missed detections.
Remarkably, the TTC system significantly improves the offline 3D detectors for test-time performance. Specifically, with test-time rectification, TTC improves offline monocular [67], multi-view [57], and BEV detectors [62] by 5.0%, 12.7%, and 12.1% EDS  [^2] EDS is a class-agnostic version of nuScenes Detection Score (NDS) [1], treating all objects as class-agnostic entities and ignoring the velocity and attribute to evaluate out-of-distribution 3D detection. , respectively. Second, on challenging scenarios, the TTC system exhibits even more substantial gains, with improvements of 14.4%, 21.6%, 13.6%, and 4.7% EDS on tasks like distant 3D detection and vehicle-focused detection with limited annotations, zero-shot extensions, as well as scenarios with domain shifts, respectively.
This comprehensive evaluation highlights the versatility and adaptability of TTC, which can effectively rectify online errors and maintain robust 3D detection even under limited data, category shifts, and environmental changes. We hope the introduction of TTC will inspire the community to further explore the online rectification approach in autonomous driving systems, a crucial technology that can enhance the safety and reliability of safety-critical applications.
2 Related Work
In this paper, we introduce the task of online 3D detection, which focuses on the instant rectification of errors during the online testing phase of 3D detectors. The primary goal is to enable the continuous detection of objects missed by offline-trained 3D detectors, without requiring additional training. This work is closely related to several areas, including 3D detection, tracking, test-time adaption and interactive vision models. Here, we discuss 3D detection and tracking, while leave others in the supplementary material Appendix B.
3D Object Detection. 3D object detection is a cornerstone task in autonomous driving. Traditional 3D detectors primarily focus on the performance of pre-trained models [50, 25, 65, 44, 27]. Once deployed, the model’s output is fixed during inference, with no ability to intervene in real time to correct missed detections. In recent years, some works have explored instruction-based detection systems, offering interfaces that accept human prompts [4, 7, 66, 15, 70, 2, 55]. However, these systems typically rely on boxes, points, or language as interfaces. Boxes and points struggle to handle temporal data [7, 66, 15], while language prompts often depend on LLMs [4, 70, 2, 55], leading to higher latency and potential ambiguities. In contrast, online 3D detection focuses on real-time error correction during sequential detection. By utilizing visual prompts, we enable accurate correction of target objects in streaming data instantly.
Target Object Tracking. Tracking is another direction closely related to online 3D detection [28, 41, 42, 68, 18]. Since our approach uses visual prompts to identify targets, it closely resembles single object tracking (SOT) in terms of its objectives [61, 5, 10, 11]. While current SOT methods based on images are limited to generating 2D detection outputs and presume the object query as the initial frame of the ongoing video, our TTC system stands out. It can directly identify and track the associated 3D bounding boxes using visual prompts, even when these prompts originate from varied sources, scenes, and zones with distinct styles.
3 Test-time Correction with Human Feedback
Figure 4: Overall Framework. (Left:) The TTC system centers on a TTC-3D Detector which utilizes visual prompts $\mathcal{P}_{v}$ from the visual prompt buffer for test-time error rectification. (Right:) The TTC-3D Detector can be based on any traditional detector (BEV or monocular). It supports 3D detection from any combination of four prompts, i.e., object $\mathcal{P}_{o}$, box $\mathcal{P}_{b}$, point $\mathcal{P}_{p}$, and novel visual prompts $\mathcal{P}_{v}$, arbitrary views of target objects across scenarios and timestamps.
In this section, we elaborate on our TTC, an online test-time error rectification system for 3D detection to detect and track missed objects during on-road inference with the guidance of human feedback. We start with an overview in Section 3.1, then delve into OA module and the visual prompt buffer in Section 3.2, and Section 3.3, respectively.
3.1 Overview
We convert online human feedback, i.e., clicks, boxes, or images from the Internet, into a uniform representation called “visual prompts”, image descriptions of target objects. Such image-based descriptions can cover arbitrary views of objects, including pictures taken from diverse zones, weather, timestamps, or even from out-of-domain sources such as stylized Internet images. Upon visual prompts, the TTC system, a recurrent framework, is designed to engage in sustained interaction with human users, continuously learning to detect and track new objects. As Figure 4 (left) shows, it comprises two key components: 1) TTC-3D Detector, any 3D detector equipped with OA module for in-context 3D detection and tracking via visual prompts, and 2) an extendable visual prompt buffer storing visual prompts of all previously missed objects, enabling continuous online error rectification.
During online inference after being deployed on cars, whenever an error occurs, i.e., miss an object, users can add the unrecognized object to visual prompt buffer by clicking on it in the image. The model then detects the corresponding 2D boxes based on the user-provided clicking prompt and updates the visual prompt buffer with the associated image patch. In subsequent frames, TTC-3D Detector leverages the stored visual prompts to detect and track previously missed 3D objects. This enables instant error correction and continuous improvement of 3D detection during online operation.
3.2 Online Adapter (OA)
OA module is designed as a plug-and-play module. It receives human prompts and transforms them into queries that can be processed by traditional detectors. When trained with OA module, traditional detectors can understand human prompts to produce detection results, enabling online error correction without the need for parameter updates.
Prompt-driven Query Generating. As shown in Figure 4 (right), OA module is flexible to handle four prompts in different forms: object query prompts for traditional offline 3D detection, box and point prompts for collecting test-time feedback, and visual prompts for consistent error correction. Specifically, these prompts are processed as follows:
  • For object prompts $\mathcal{P}_{o}$, the OA module generates a set of learnable embeddings as queries, akin to traditional 3D detectors, which are updated during training as demonstrated in previous works [27, 56, 67].
  • For box $\mathcal{P}_{b}$ and point $\mathcal{P}_{p}$ prompts, the OA module encodes them with their location and shape, representing them as Fourier features [51].
  • For visual prompts $\mathcal{P}_{v}$, the OA module first extracts the visual prompt features $\mathcal{Z}_{v}$ by an encoder [17], then localizes their corresponding objects within the input images $\mathcal{X}_{v}$, and finally adds their features with the Fourier positional encoding as subsequent inputs:
    $\mathcal{P}_{v}=\texttt{FourierPE}(\texttt{Align}(\mathcal{Z}_{v},\mathcal{Z}))+\mathcal{Z}_{v},$ (1)
    where $\mathcal{Z}$ means image features of image input, $\texttt{FourierPE}(\cdot)$ is the Fourier positional encoding and the $\texttt{Align}(\cdot)$ operation means “Visual Prompt Alignment”, the process of inferring the 2D position in the current frame of the visual prompt.
Visual Prompt Alignment. We perform the Align operation to localize target objects in the input images, by visual prompts. To handle flexible visual prompts, which can be image descriptions of objects in any view, scene, style, or timestamp, we employ contrastive mechanisms [59, 16] as the key design to retrieve target objects at different styles. This allows the module to detect the target objects effectively, even when they exhibit diverse visual styles and appearances.
Specifically, we use two multi-layer perceptrons (MLPs) to first align the channels of image features $\mathcal{Z}$ and visual prompt features $\mathcal{Z}_{v}$. We then compute the dot product between the aligned feature maps to obtain a similarity map. To further retrieve the coordinates of target objects from the similarity map, we multiply it with the original image features $\mathcal{Z}$, and apply another MLP with channels of 2 to regress the spatial positions $\mathcal{X}_{v}$ of target objects.
Instance Ambiguity & Loss. Sometimes, visual prompts might exhibit instance ambiguity, where multiple objects in the image match the visual descriptions of prompts. For example, suppose the visual prompt describes a traffic cone and several similar-looking traffic cones present in the input images. It can be challenging to uniquely identify the specific object-of-interest.
For such cases, we design to retrieve all objects with similar identities to the visual prompt. Specifically, we modify the align operation to predict multiple spatial coordinates $\mathcal{X}_{v}=\{\mathcal{X}^{(i)}_{v}\},i\in\{1,2,...,N\}$ for each visual prompt, and add the Fourier features of those $N$ coordinates to the visual prompt features to indicate visual prompts in different positions. $N$ is set to 4 in our implementation.
For similarity supervision, we generate binary segmentation labels based on the ground-truth 2D boxes. Focal loss [30] and Dice loss [37] are used for optimization. To supervise the visual prompt localization $\mathcal{X}_{v}$, we use the Smooth-$l$1 loss with the target as the center coordinates of ground-truth 2D bounding boxes. For dealing with instance ambiguity, we refer to SAM [24] and only backpropagate the sample with the minimum localization loss during each training iteration. For more details, please refer to the supplementary material Section C.2.
Model Design. The overall mechanism of TTC-3D Detector is depicted in Figure 4 (right). Based on any traditional offline-trained 3D detector, BEV detector, or monocular detector, we integrate OA module and train it to be promptable. Specifically, OA module takes features extracted by the image encoder of the corresponding 3D detector, along with various forms of prompts as inputs. It encodes these prompts and generates a series of queries, represented as $\mathcal{P}=\{\mathcal{P}_{o};<\mathcal{P}_{b},\mathcal{P}_{p},\mathcal{P}_{v}>\}$, where <>expectation<...>< … > denotes an arbitrary combination of different prompts. These queries are then fed into the transformer decoder of the 3D detector to output 3D boxes following human feedback.
3.3 Visual Prompt Buffer
Visual prompt buffer is a queue that stores user-provided visual prompts of missed objects during online inference. In the default setting, these visual prompts are the 2D output regions from the model in previous frames, based on the user’s point or box prompts at that time. The flexibility of the TTC system also allows users to freely select visual prompts, either image contents from the current scene or customized objects from the Internet. This versatility makes the TTC system applicable to a wide spectrum of scenarios.
Dequeue. To prevent the buffer from growing indefinitely, we design a “dequeue” mechanism. We filter out visual prompts with low confidence, which likely no longer appear in the scene. We also use the intersection-over-union (IoU) between predictions to identify and remove redundant visual prompts. This dynamic update and maintenance ensure a balance of TTC system between latency and accuracy by adaptively incorporating and pruning feedback online.
4 Experiments
We now proceed to evaluate our TTC system from two key aspects:
  • How does TTC perform with various adapted offline 3D detectors for instant online error rectification?
  • How does TTC enhance offline-trained 3D detectors when faced with out-of-training-distribution scenarios?
4.1 Experimental Setup
Dataset. Experiments are done on nuScenes dataset [1] with 1,000 autonomous driving sequences, one of the most popular datasets for autonomous driving research.
Tasks. Experiments are conducted in two aspects to answer the questions above. The first experiment aims to test the TTC system in boosting the performance of offline-trained 3D detectors via online error rectification. We conduct this verification by applying TTC system to various established offline 3D detectors [67, 27, 62, 32, 54, 31, 57].
Then, we validate the TTC system to correct detection errors in out-of-training-distribution scenarios. To conduct a thorough quantitative analysis of this aspect, we established four tasks under different settings: (a.) discarding 80% labels of distant objects farther than 30m during training, and test the error corrections for detecting distant objects; (b.) discarding 80% labels of instances labeled as vehicle, including “car”, “truck”, “C.V.”, “bus”, and “trailer”, and test the error correction in vehicles; (c.) discarding all annotations of class “truck” and “bus”, and test the zero-shot ability of TTC on those discarded classes; (d.) discarding all training data of the scenario of “Nighttime” and “Rainy”, and test the improvements of TTC on scenarios with domain gap. For these tasks, we base TTC system on the monocular algorithm, MonoDETR [67], and test it under each setting separately, as the monocular setting demonstrates the most general applicability. Based on TTC-MonoDETR, we also present qualitative results to show the potential of TTC to address corner cases in challenging real-world scenarios through human feedback during test time. Through these experiments, we demonstrate TTC as an effective system to adapt offline 3D detectors to challenging scenarios.
Entity Detection Score.
As we focus on enabling existing 3D detectors to rectify missing detections, both in-distribution and out-of-distribution, we adopt a class-agnostic setting. We remove all class annotations of 3D objects during training, treating all objects as entities [24, 45]. For evaluation, we use the Entity Detection Score (EDS), a class-agnostic version of the nuScenes Detection Score (NDS) that emphasizes recall. A detailed analysis of the rationale behind EDS can be found in Appendix Section C.3.
Implementation Details. We implement our method based on MMDetection3D codebase [8], and conduct all experiments on a server with 8$\times$ A100 GPUs. In OA module, we use a ResNet18 [17] to extract visual prompt features. All visual prompts are resized to $224\times 224$ before being sent to OA module. For training, we use AdamW [23, 36] with a batch size of 16. We initialize the learning rate as 2e-4, adjusted by the cosine annealing policy. When training point and box prompts, we simulate user inputs with noise by adding perturbations to the ground truth. To ensure the visual prompts are robust across multiple scenes, timestamps, and styles, we choose visual prompts of target objects not only from the image patches of the current frame but also randomly from previous and future frames within a range of ±$\pm 5$. Flip operations are used as data augmentations. During testing, we mimic user intervention. We simulate the user’s online missing feedback by comparing the distance between ground truth and detected 3D boxes. Ground-truths without any detection within 2m are considered missed and added into the prompt buffer. We remove redundant predictions by non-max-suppression (NMS) with the IoU threshold of 0.5 and set the classification confidence threshold at 0.3.
Table 1: Effect of TTC system on various 3D detectors. TTC system effectively improves the test-time performance of offline-trained detectors during online inference without any extra training.
Method Backbone Type mAP (%) $\uparrow$ EDS (%) $\uparrow$
MonoDETR [67] R101 Monocular 37.9 38.2
TTC-MonoDETR 42.6 (+5.7) 43.2 (+5.0)
MV2D [57] R50 Multiview 38.9 38.3
TTC-MV2D 51.0 (+12.1) 51.0 (+12.7)
Sparse4Dv2 [31] R50 Multiview + Temporal 38.8 37.6
TTC-Sparse4Dv2 53.5 (+14.7) 52.1 (+14.5)
BEVFormer [27] R101 BEV + Temporal 36.5 35.8
TTC-BEVFormer 47.8 (+11.3) 47.2 (+11.4)
BEVFormerV2-t8 [62] R50 BEV +Temporal 39.6 38.9
TTC-BEVFormerV2-t8 51.6 (+12.0) 51.0 (+12.1)
RayDN [32] R50 BEV + Temporal 39.6 38.7
TTC-RayDN 51.7 (+12.1) 50.8 (+12.1)
StreamPETR [54] V2-99 BEV + Temporal 39.7 39.1
TTC-StreamPETR 52.3 (+12.6) 51.5 (+12.4)
Table 2: Experiments on out-of-training-distribution scenarios. TTC system achieves substantial gains with limited or even no labels under different challenging test cases.
(a) Long-range rectification. Effect of TTC system in detecting distant objects with 20% annotations.
Model Setting All (0m-Inf) Dist. (30m-Inf)
mAP (%) EDS (%) mAP (%) EDS (%)
Point 41.6 40.8 17.8 19.4
Box 44.3 43.7 19.2 21.7
Visual 42.6 42.0 18.3 21.6
MonoDETR 31.4 31.6 0.0 0.0
TTC-MonoDETR 40.2 40.1 11.0 14.4
$\Delta$ +8.8 +8.5 +11.0 +14.4
(b) Vehicle-focused rectification. Effect of TTC system on vehicle objects with 20% annotations.
Model Setting All Vehicle
mAP (%) EDS (%) mAP (%) EDS (%)
Point 38.0 36.9 33.5 36.6
Box 41.2 40.0 36.7 39.6
Visual 39.2 38.4 34.6 38.5
MonoDETR 17.6 16.1 2.4 7.3
TTC-MonoDETR 29.0 27.2 23.2 28.9
$\Delta$ +11.4 +11.1 +20.8 +21.6
(c) Novel object rectification. Effect of TTC system on objects of novel classes unseen in the training set.
Model Setting All Unseen
mAP (%) EDS (%) mAP (%) EDS (%)
Point 35.2 35.6 11.9 15.0
Box 38.5 38.8 14.7 16.9
Visual 35.4 36.0 11.2 15.6
MonoDETR 28.3 29.6 0.0 0.0
TTC-MonoDETR 34.8 36.1 8.5 13.6
$\Delta$ +6.5 +6.5 +8.5 +13.6
(d) Domain shift rectification. Effect of TTC system on objects in scenarios with domain gap.
Model Setting All Rain & Night
mAP (%) EDS (%) mAP (%) EDS (%)
Point 39.0 38.2 30.5 30.7
Box 42.6 41.5 33.4 33.6
Visual 39.8 39.4 29.4 30.8
MonoDETR 34.5 34.7 25.2 26.6
TTC-MonoDETR 39.9 40.0 29.7 31.3
$\Delta$ +5.4 +5.3 +4.5 +4.7
Figure 5: Qualitative visualization of real-world scenes (collected from YouTube). We visualize the zero-shot 3D detection results in a real-world scenario. In this case, the prompt buffer contains a visual prompt of a deer. Higher responses from the visual prompt alignment are highlighted by brighter colors. As shown, although trained solely on nuScenes, TTC system can still accurately localize “unseen” objects in the input image. Best viewed in color.
4.2 Main Result
Effectiveness of TTC in Test-time Error Correction. Test-time error correction ability without re-training is the core capability of TTC system. We verify this by incorporating traditional offline-trained 3D detectors into TTC system and compare the performance without test-time parameter updates. For thorough verification, we select various offline-trained 3D detectors, including monocular [67], multi-view [57, 31], and BEV ones [27, 62, 54, 32]. As shown in Table 1, the TTC system substantially improves the test-time performance of offline-trained 3D detectors, e.g., 11.4% and 12.4% EDS improvements on BEVFormer and StreamPETR, without requiring additional training during inference. These demonstrate the effectiveness of the TTC system in instantly correcting test-time errors during online inference.
Effectiveness of TTC in Out-of-Training Scenarios. Table 2 presents results to validate the TTC system in challenging and extreme scenarios. In these experiments, we base TTC system on MonoDETR for its simple monocular setting. “Point”, “Box”, and “Visual” in Table 2 are TTC-MonoDETR using point, box, or image patch of target objects in the input image as input prompts. These are performance upper bounds as they receive human feedback in every frame.
Table 2(a) evaluates the effectiveness of TTC system in rectifying detection errors on distant objects (beyond 30m). During training, 80% of the far-away annotations are removed, resulting in a 3D detector with poor long-range performance (0.0% mAP and EDS). Powered by the TTC system, the performance on these distant objects is instantly improved to 11.0% mAP and 14.4% EDS, without any extra training. The experiment is then extended to all vehicles in the nuScenes dataset, as shown in Table 2(b). The TTC achieves impressive performance gains of 20.8% mAP and 21.6% EDS.
Furthermore, in Table 2(c) and Table 2(d), we evaluate the TTC system in two challenging scenarios: encountering unseen objects not present in the training data; handling domain shifts, such as transitioning from sunny to rainy or nighttime conditions. In Table 2(c), the TTC system shows effectiveness in successfully detecting novel objects not labeled in the training set, achieving 8.5% mAP and 13.6% EDS on these novel objects, significantly improving the offline-trained baseline with 0.0% mAP and EDS. In Table 2(d), we find TTC also works well for online error rectification when driving into scenarios with domain shifts, providing 4.5% mAP and 4.7% EDS improvements.
Regarding qualitative results, Figure 5 further illustrates a case of the zero-shot capability in the real-world. Despite being trained solely on the nuScenes dataset, TTC-MonoDETR can detect a Deer using a visual prompt of another deer from a different viewpoint, which is extremely challenging for traditional offline detectors. More qualitative examples can be found in supplementary material Appendix E.
These experiments demonstrate TTC as an effective and versatile system for instant error correction, excelling at handling missing distant objects, unseen object categories, and domain shifts, without requiring any additional training. The superior performance of TTC system in these challenging real-world scenarios highlights its potential to enable robust and adaptable 3D object detection systems.
4.3 Robustness of Visual Prompts
As a crucial component of the continuous test-time error correction system, we conduct a series of ablation studies on visual prompts in this section. We base TTC system on MonoDETR and validate its robustness concerning prompts obtained from the Internet or those from different moments.
Figure 6: Visual prompt examples derived from the Internet. We select 12 cars and 6 buses with different views and colors from websites to cover the distribution of object appearances in the nuScenes dataset comprehensively.
Table 3: Results with web-derived visual prompts. TTC-MonoDETR indicates the model with frozen prompt buffer containing pre-assigned visual prompts derived from the Internet. TTC system can still significantly improve traditional 3D detectors even using web prompts in scenarios with very limited labeled data.
Category Method mAP (%) $\uparrow$ EDS (%) $\uparrow$
Car MonoDETR 3.2 9.1
TTC-MonoDETR 20.9 (+17.7) 27.5 (+18.4)
Bus MonoDETR 0.4 4.4
TTC-MonoDETR 17.4 (+17.0) 20.6 (+16.2)
Figure 7: Experiments on visual prompts from arbitrary temporal frames. TTC-3D Detectors can effectively locate and detect target objects in future frames, using its image patch at Frame #0.
Robustness over Web-derived Visual Prompts. Visual prompts can be arbitrary imagery views of target objects and can be from any image source. For example, we can use images sourced from the Internet as visual prompts. We fix the prompt buffer with visual prompts from the Internet during inference, and assess the effectiveness of handling visual prompts with diverse styles. We employ the model from Table 2(b) (TTC-MonoDETR trained with 20% labels of vehicles), and select 12 car and 6 bus images from the Internet, which resemble those in nuScenes dataset, as visual prompts for online correction. We show them in Figure 6. As listed in Table 3, TTC system demonstrates strong online correction capabilities though with limited annotations. Even with prompts sourced from the Internet with various styles and poses, TTC system still improves the offline-trained baseline by 18.4% EDS on “Car” objects under this extremely challenging setting. This further underscores the robustness of TTC system and highlights its potential to address the long-tail challenges in real-world scenarios.
Robustness over Arbitrary Visual Prompt Views. We also investigate the capability of TTC system in handling visual prompts of target objects across scenes and times. In Figure 7, we study whether our system can successfully associate objects with its visual prompts from arbitrary frames. Specifically, for each video clip of the nuScenes validation set, we use the image patch of target objects in the first frame as visual prompts, then detect and track target objects in subsequent frames, and compute the recall rate for evaluation with arbitrary views. Figure 7 presents the results of TTC-MonoDETR, showing that despite significant differences in viewpoints and object poses between frames, the recall rate does not drop dramatically as the ego vehicle moves. This highlights the robustness of the TTC detectors in handling visual prompts with arbitrary views across scenes and times, indicating that the TTC system can effectively process human feedback that may involve temporal delays.
4.4 More Ablation Studies
Due to limited space, we refer readers to Appendix D of the appendix for extensive ablation studies. These studies include: (1.) the ablations on core components in visual prompt alignment in Section D.1 and Section D.2, (2.) the evaluation of visual prompt alignment’s impact on ensuring instance awareness in Section Section D.3, (3.) the examination of statistics regarding the visual prompt buffer during online inference in Section D.4, (4.) the assessment of visual prompt robustness against user perturbations and reduced user feedback collection frequency in Section D.5 and Section D.6, (5.) the comparison highlighting the superiority of our OA module over the combination of 2D single-object trackers with 3D heads for generating 3D tracking in Section Section D.7, and more.
5 Conclusion
In this paper, we introduce the TTC system. It equips existing 3D detectors with the ability to make test-time error corrections. The core component is the OA module, which enables offline-trained 3D detectors with the ability to leverage visual prompts for continuously detecting and tracking previously missing 3D objects. By updating the visual prompt buffer, TTC system enables continuous error rectification online without any training. To conclude, TTC provides a more reliable online 3D perception system, allowing seamless transfer of offline-trained 3D detectors to new autonomous driving deployments. We hope this work will inspire the development of online correction systems.
Limitations and Future Work. As the first work on online 3D detection, this study has several limitations, which might open up many potential research directions.
For performance, the generalization ability is constrained by the model size and data volume, and future work will aim to enhance this by scaling up both the data and parameters.
For model design, while allowing user interaction provides reliability, it introduces potential risks. Requiring frequent human feedback may be a challenge in current L3 autonomous driving paradigms. Though reducing the frequency of feedback can alleviate this pressure without compromising performance (See appendix Section D.6), there is room for improvement. Future work could integrate robust tracking methods or leverage advanced hardware, such as eye-tracking, to simplify the feedback collection.
For objectives, this work focuses on addressing missed detections. It does not tackle the redundancy issues. Future work could explore addressing false positive issues through TTC system. Additionally, to enable zero-shot detection, this work adopts a class-agnostic approach. In the future, methods combining language models could be explored, ensuring both low latency and open-world category detection ability. Finally, this study only explores error correction within the visual domain; extending this approach to other sensors, such as LiDAR, will be an area for further investigation.
Acknowledgement
OpenDriveLab is the autonomous driving team affiliated with Shanghai AI Lab. This work was supported by National Key R&D Program of China (2022ZD0160104), NSFC (62206172), and Shanghai Committee of Science and Technology (23YF1462000). We thank team members from OpenDriveLab for valuable feedback along the project. Special thanks to Chengen Xie for providing data support, and Jiahui Fu for sharing insights on MV2D.
Appendix A Discussions
For a better understanding of our work, we supplement intuitive questions that one might raise. Note that the following list does not indicate whether the manuscript was submitted to a previous venue or not.
Q1: What is the relationship between the online 3D detection and the offline data-loop progress?
We emphasize, in this paper, we do not propose the online system to replace the traditional offline data loop. As illustrated in Fig. 2 of the main paper, these two systems address different aspects of autonomous driving. The offline system remains crucial for enhancing the capabilities of the base perception model through development; while our online TTC system further enables the deployed frozen model in vehicles to promptly rectify dangerous driving behaviors caused by unrecognized objects on the road. With improved offline-trained detectors, the TTC can effectively correct more online errors during test-time inference without re-training, as detailed in Table 1 of the main paper.
Q2: What are the main technical novelty and advantages of the proposed TTC system over previous instruction-based 3D detectors?
The advantages lie in the design of visual prompts, the visual descriptions of target objects with diverse sources, styles, poses, and timestamps. While existing instruction-based 3D detection methods typically utilize text, boxes, or clicks as prompts.
Compared to text prompts, visual prompts provide a more natural and accurate description of the target object. In contrast, verbal descriptions can be ambiguous to convey instance-level features, leading to an inaccurate understanding of the missing objects. Second, text promptable models are often combined with LLMs with high latency and are thus unavailable for autonomous driving deployment, as discussed in Section D.8.
Box and point prompts are less convenient than visual prompts when dealing with stream data. If missing occurs, these single-frame prompts require users to provide feedback at every frame, which is unfeasible in real-world applications. Compared to box and point prompts, visual prompts are robust across different scenes and timestamps, one single-frame visual prompt is enough for detecting and tracking in later frames. Furthermore, visual prompts enable 3D detection with pre-defined visual descriptions of target objects, regardless of the sources, styles, poses, etc, as discussed in Section 4.3.
The introduction of novel visual prompts enables real-time, accurate, and continuous error correction of streaming inputs with ”one-click” feedback.
Q3: What is the relationship between the TTC system and existing 3D perception tasks?
The TTC system relates to several 3D perception tasks, including 3D object detection, zero/few-shot detection, domain adaptation, single object tracking, and continual learning.
Compared to standard 3D object detection, the primary focus of the TTC system is on enabling instant online error correction rather than optimizing the offline detection performance of the base 3D detector. In contrast to traditional few-shot, one-shot, or domain adaptation approaches, the TTC system does not require 3D annotations for new objects or any model retraining, yet can still provide reasonable 3D bounding box estimates for out-of-distribution objects. Relative to image-based single object tracking, TTC is not merely limited to generating 2D detection outputs and assuming the object query as the initial frame of an ongoing video. Instead, it is capable of performing 3D tracking based on visual descriptions of target objects from any scene or timestamp, leveraging a diverse set of visual prompts.
In summary, the TTC system represents a more flexible and comprehensive 3D object detection framework, combining the strengths of zero-shot detection, handling out-of-distribution objects, and utilizing diverse visual prompts beyond the current scene context.
Q4: Why choose 3D detection as the experimental scenarios of TTC? Could the proposed framework be extended to 2D detection or other vision tasks?
TTC represents a general idea to equip deployed systems with the capability of online error rectification, making them more versatile, adaptive, and reliable. This idea can be readily extended to other vision tasks, such as 2D detection, for rapid adaptation of pre-trained models to novel scenarios.
The choice of 3D detection for autonomous driving as the experimental scenario is motivated by the paramount importance of safety for deployed self-driving systems. Without an online correction method, mistakes made by the offline model pose significant safety risks for on-road autonomy.
Therefore, we select the autonomous driving domain as the testbed for the TTC framework, given the critical need for a robust, adaptive online error correction system to ensure the reliability of these safety-critical applications. We mark the extension of TTC to other vision tasks as future works.
Q5: What are potential applications and future directions of TTC?
We believe that, visual prompts, as the core design element of the TTC system, represent a more natural and intuitive query modality for the image domain. This approach has significant research potential and application prospects in the field of 3D perception and beyond.
For example, visual prompts enable rapid customization of the tracking targets, beyond the pre-defined object classes. Second, the visual prompt-based framework facilitates online continual learning for 3D perception systems, adapting to evolving environments. Then, visual prompts can be applied in the V2X domain to enable swift error rectification across diverse operational scenarios. Visual prompts can also be deployed to assist in the auto-labeling process of target objects. Furthermore, by combining visual prompts with natural language prompts, we can obtain more precise descriptions and behavioral control for online perception systems.
The diverse applications outlined above demonstrate the promise of visual prompts as a versatile approach. As showcased in this work, the visual prompt-based framework opens up new possibilities for online perception systems, not only in autonomous driving but also in a broader range of domains.
Appendix B More Related Work
B.1 Interactive Vision Models.
Interactive vision models are designed for tasks based on user inputs. As one of the fundamental tasks in computer vision, they are extensively researched with numerous breakthroughs [26, 6, 33, 60, 14]. Particularly, the advent of the Segment Anything Model (SAM) [24] has sparked a surge of progress, with applications spanning reconstruction [49], detection [47, 63], segmentation [71], image editing [13], and more. Compared to existing models, which typically rely on prompts such as clicks, boxes, or scribbles, in this work, we study visual prompts, a new prompt referring to the actual images of objects with arbitrary poses and styles. Visual prompts enable continuous detection and tracking of target objects, facilitating the immediate correction of failed detections at test time. This represents a departure from traditional prompt types, offering more natural and dynamic interactions with the visual content.
B.2 Test-time Adaptation
Test-time adaptation aims to improve model performance on test data by adapting the model to test samples, even in the presence of data shifts [69, 38, 39, 48, 53, 34]. However, this area remains relatively unexplored in the context of 3D detection. Recent work, such as MonoTTA [29], has explored domain adaptation for 3D detection by fine-tuning models on batch streaming data under different weather conditions. In contrast, online 3D detection does not modify the original model parameters. Instead, it leverages human feedback to achieve instance-level error correction, thereby enabling domain adaptation without compromising the model’s performance on the original domain.
Appendix C Implementation Details
C.1 Details of TTC System
Figure 8: Details of the TTC system.
This section elaborates on the detailed workflow of the TTC system. As shown in Figure 8, during online inference, once the TTC-3D Detector fails to recognize an object, users can click on the missing object within the image. Based on the user-provided click, the TTC-3D Detector identifies the corresponding 2D boxes, crops the relevant areas to obtain visual prompt patches, and subsequently updates the visual prompt buffer. In later frames of inference, the TTC-3D Detector applies these stored visual prompts to continuously detect and track previously missing 3D objects, achieving instant error correction during test-time and constantly enhancing the offline-trained 3D detectors after being deployed.
C.2 Visual Prompt Alignment
Figure 9: Concrete mechanism of visual prompt alignment. This figure illustrates monocular input. When multi-view images are employed, this alignment operation flattens the different views and still generates $N$ peak candidate positions.
We now delve into the implementation details of visual prompt alignment. As described in Figure 9, given the visual prompt $\mathcal{P}_{v}$, we first extract the visual prompt features $\mathcal{Z}_{v}$ using a lightweight encoder, e.g., ResNet18 [17] in our implementation. Then, we use two separate 2-layer perceptrons, each with 128 and 64 output channels, to align the channel dimensions of the prompt features and the input image features. This results in the aligned feature maps $\hat{\mathcal{Z}_{v}}\in\mathbb{R}^{M\times 64}$ and $\hat{\mathcal{Z}}\in\mathbb{R}^{KHW\times 64}$, where $M$ is the number of visual prompts, $K$ is the number of image views, $H$ and $W$ are the spatial dimensions of image features. Finally, we compute the cosine similarity ($\odot$ in Figure 9) between $\hat{\mathcal{Z}}_{v}$ and $\hat{\mathcal{Z}}$ to obtain the similarity map $\mathcal{S}\in\mathbb{R}^{M\times KHW}$, which encodes the alignments between the visual prompts and image features.
To solve the instance ambiguity issue, we propose to predict multiple spatial coordinates of different peak responses in the similarity map (highlighted by “orange box” in Figure 9). Specifically, we first multiply the image features $\mathcal{Z}$ with similarity map $\mathcal{S}$ ($\bigotimes$ in Figure 9), and then use a 2-layer perceptron with $N\times 2$ output channels to regress spatial coordinates of the $N$ peak responses. During training, we only backpropagate the localization loss for the positions that have the minimum loss with respect to the ground truth. In our implementation, $N$ equals 4. This multi-instance retrieval approach allows the TTC to handle cases where the visual prompt matches multiple candidate objects in the input image, improving the robustness of the online error correction.
C.3 Details of Entity Detection Score (EDS)
Table 4: Feasibility Analysis of the EDS Metric.
Method mAP (%) w.o. Rec (%) + Rec (%)
BEVFormer 36.5 49.1 35.8
TTC-BEVFormer 47.8 54.6 47.2
BEVFormer-V2 39.6 49.2 38.9
TTC-BEVFormer-V2 51.6 58.5 51.0
EDS is a class-agnostic version of nuScenes Detection Score (NDS), but further prioritizes the localization quality of target objects. Inspired by [64], we improve the original NDS computation by multiplying the recall rate and the mean True Positive metrics, $\mathbb{TP}$, as well, as illustrated below:
$\text{EDS}\!=\!\frac{1}{6}[3\text{mAP}\!+\!\text{Recall}\!\times\!\sum_{\text{mTP}\in\mathbb{TP}}(1-\text{min}(1,\text{mTP}))],$ (2)
The intuition behind this is simple. The larger the recall rate is, the more predictions are involved in the statistics of mTP. Compared to simply setting a recall threshold [1], the multiplication adjusts the weight of mTP to EDS according to its comprehensiveness and thus brings a more informative quantitative result. In Table 4, we analyze the effectiveness of this metric. The comparisons between multiplying recall rate or not on various TTC-3D Detectors show that EDS, incorporating recall into its calculation, does not alter the original overall trend. Furthermore, it effectively highlights the superiority of detecting missed objects, demonstrating its validity.
Appendix D Ablation Studies
In this section, we conduct a series of ablation studies. We base our TTC system on MonoDETR for its simple and efficient monocular setting except for experiments in Section D.6, Section D.7 and Section D.9.
D.1 Effect of Components in Visual Prompt Alignment.
Table 5: Effect of visual prompt alignment.
Sim. Loss Loc. Loss mAP (%) EDS (%)
- - 32.6 31.9
$\surd$ - 39.4 38.5
$\surd$ $\surd$ 43.3 42.8
Visual prompt alignment aims to localize objects via visual prompts in input images. We now evaluate its components. Table 5 presents experiments to verify the core components of this alignment, including similarity loss (Focal and Dice loss) and localization loss (supervising visual prompt localization, $\mathcal{X}_{v}$). While the alignment can be implicitly learned by attention mechanisms in the transformer decoder, incorporating explicit similarity supervision brings improvements of 6.8% mAP and 6.6% EDS. Further utilizing the position loss and one-to-$N$ mapping (to address instance ambiguity) boosts the performance to 43.3% mAP and 42.8% EDS. These results prove this alignment operation is a critical component enabling the TTC detectors to effectively detect target objects via visual prompts.
D.2 Instance Ambiguity in Visual Prompt Alignment
Table 6: Effect of the number of predicted positions $N$ of visual prompt alignment.
No. of Position Predictions mAP (%) $\uparrow$ Recall (%) $\uparrow$
1 39.9 62.1
4 43.3 69.1
8 43.0 69.4
To solve the instance ambiguity issue, we propose to regress $N$ positions of each visual prompt when performing the visual prompt alignment. This retrieves all objects with similar visual contents. In this study, we validate the effectiveness of this design by conducting ablation studies on the number of $N$. As listed in Table 6, when the number of position prediction $N$ equals 1, which means a one-to-one mapping for each visual prompt, the mAP and recall rate are 39.9% and 62.1%, respectively. Then, if we increase the $N$ to 4, effectively a one-to-four mapping, we obtain an mAP of 43.3% and a recall rate of 69.1%. This represents a 7% improvement in the recall rate, demonstrating that instance ambiguity is an important challenge in visual prompt-based detection, and the proposed one-to-$N$ mapping solution effectively addresses this issue.
D.3 Effect of Visual Prompt Alignment on Instance Awareness
Figure 10: Size of visual prompt buffer during the video stream. Visual prompt buffer stores the missed objects during online inference to rectify test-time errors of deployed 3D detectors. It can adaptively manage the stored prompts and thus maintain the balance between latency and accuracy.
Figure 11: Visual prompt showcase for Table 7.
Table 7: Effect of visual prompt alignment on instance awareness. Group (a.) resembles “car” objects in nuScenes validation set with diverse types and colors; while group (b.) maintains black sedans only. The performance gap demonstrates that alignment learns the instance awareness.
Group mAP (%) $\uparrow$ EDS (%) $\uparrow$
a. 20.9 27.5
b. 13.8 18.9
Despite demonstrating that visual prompt alignment can enhance system performance, we remain uncertain whether the alignment can distinguish different objects based on visual prompts for instance-level matching. To investigate, we conduct an additional experiment, based on a similar setting with Table 3 of the main paper [^3] For reference, we employ the TTC-MonoDETR trained on 20% vehicle annotations and freeze the prompt buffer with predefined web prompts when inference., but fix the visual prompt buffer with images of black sedans solely (Figure 11). As shown in Table 7, as “Car” objects in nuScenes contain various types and colors, solely using black sedans as visual prompts leads to an 8.6% EDS drop. This underscores that the alignment operation effectively differentiates objects based on visual prompts, achieving instance-level matching and detection.
D.4 Statistics of Visual Prompt Buffer during Online Inference.
We design the visual prompt buffer to store missed objects during inference with video stream and introduce a “dequeue” mechanism to prevent the buffer from growing indefinitely. In this ablation study, we analyze the dynamic buffer size, as well as the number of enqueued and dequeued in each frame, to illustrate the behavior of visual prompt buffer during the online operation of TTC.
As shown in Figure 10, the visual prompt buffer exhibits three distinct behaviors during online inference in each nuScenes video clip: increasing, steady, and decreasing. In initial frames, many traffic cones are queued into the buffer due to the poor performance of deployed offline 3D detectors on cone objects. The buffer size thus grows quickly in initial frames to store visual prompts of missed objects for online rectification (Frames #0 to #4). The buffer size then stabilizes as the online detector consistently detects and tracks all objects of interest (Frames #4 to #20). Further, as the ego vehicle drives out of the scene, many previously enqueued objects no longer exist and are thus removed from the buffer automatically (Frames #20 to #40). The buffer finally becomes empty as the scene changes.
This demonstrates the effectiveness of visual prompt buffer, which consistently stores missed objects during online inference and corrects online errors. This dynamic behavior, exhibiting increasing, steady, and decreasing phases, highlights itself to manage stored visual prompts for robust 3D object detection performance throughout the online inference with the balance between latency and accuracy.
D.5 Visual Prompts with User Pertubations
Figure 12: Robustness against feedback noise.
In Section 4.3, we have demonstrated the robustness of our system on visual prompts from diverse sources, styles, poses, scenes, and timestamps. Considering visual prompts during online inference are derived from user clicking, this will have positional deviations to the perfect 2D center of the target object. Thus, we design experiments with positional perturbations during test-time inference to analyze the impact. Expressly, we set the maximum translation ratios at 0%, 10%, 20%, 30%, and 40% to the ground-truth 2D centers, and input them to TTC system to evaluate the robustness against such disturbances. Figure 12 shows that the TTC-MonoDETR exhibits minor performance drops under increasing positional perturbations, demonstrating strong robustness to these disturbances and highlighting significant potential for real-world applications.
D.6 Robustness of TTC to Reduced Feedback Frequency
Table 8: mAP comparisons of different prompt collection frequencies. “N=0” means having 2D feedbacks of missed detections at every frame; “N=2” means less 2D feedbacks collected every 2 frames. TTC system demonstrates robust error correction performance even with reduced feedback frequency.
Exp. on Human Feedback collected with different frame interval $N$. N
0 2 4 6
TTC-BEVFormer 47.8 47.5 47.3 46.9
TTC-Sparse4Dv2 53.5 53.1 52.4 51.8
TTC-MV2D 51.0 50.7 50.2 49.6
Considering the practical challenges of deploying this system, where human feedback inevitably involves some delay, it is difficult to provide feedback on missed detections for every frame without additional aids. To address this, we design experiments to simulate such scenarios and investigate the robustness of visual prompts under reduced human feedback frequency. As shown in Table 8, we update the prompt buffer at varying frame intervals to mimic feedback provided at different time intervals. TTC demonstrates robust error correction performance even with reduced feedback frequency (from every frame to every 6 frames), showcasing its feasibility for real-world applications where real-time feedback may not always be practical.
D.7 Comparisons to Single Object Tracking (SOT) with 3D Detection Head
Table 9: Comparison of error correction performance with SOT + 3D detector head. The results demonstrate the superior performance of TTC in online error correction, highlighting the limitations of the baseline approach where simply combining SOT with a 3D detection head does not effectively address error correction.
Method mAP (%) $\uparrow$ EDS (%) $\uparrow$
SOT + MV2D 43.0 43.2
TTC-MV2D 51.0 51.0
In this section, we compare our TTC system with a simple implementation of online error correction that integrates SOT and 3D detection. Due to the lack of related work on single object 3D tracking, we use the ATOM algorithm for 2D tracking [12] and the MV2D detector as the 3D object detection head [57]. MV2D can leverage the 2D tracking results obtained by ATOM to initialize the query, thereby achieving the error correction ability similar to TTC system. For implementation, we record the missed objects of MV2D and store their reference images along with 2D bounding boxes in the prompt buffer, which are inputs ATOM requires. We do not perform dequeue operations to keep tracking all missing objects by ATOM.
As shown in Table 9, the TTC system, empowered by the OA module, significantly outperforms the combination of SOT and the 3D detector head in error correction performance. This result highlights that simply combining SOT with a 3D detection head does not effectively achieve online error correction, further demonstrating the effectiveness of our TTC system.
D.8 Analysis of Model Size and Latency
Table 10: Parameters and latency comparisons between LLM based 3D detectors, traditional 3D detectors, and related TTC 3D detectors.
Method LL3DA MonoDETR TTC-MonoDETR
#Params 118M 68M 80M
FPS (Hz) 0.42 11.1 9.1
We further analyze the model size and latency of TTC system, primarily comparing them with recent LLM-based promptable 3D detection methods [4, 22]. We argue that, for online promptable systems that are developed for real-world applications, latency can come from into two parts. The first one is the unavoidable delay caused by humans from observing the error to reacting to provide prompts. This is inherent to any online prompt-based approach. The other one is the delay associated with the inference speed of the system itself. As the first one cannot be controlled by the system design, we focus on the latter here.
As shown in Table 10, LLM-based promtable methods, like LL3DA [4] exhibit high latency that is inadequate for autonomous driving deployments. In contrast, our TTC system introduces only a little extra latency compared to its base detector, which meets the real-time inference requirements and thus can be applicable for online autonomous driving systems.
D.9 Comparison with Offline Fine-tuning using User Feedback
Table 11: Comparisons between the TTC-MV2D and MV2D fine-tuned with feedback 2D annotations. “N=0” means having 2D feedbacks at every frame; “N=2” means less 2D feedbacks collected every 2 frames. TTC system, though without any extra training, still outperforms the offline 2D fine-tuned MV2D, especially when annotations are limited (N=6).
Exp. on Human Feedback collected with different frame interval $N$. N
0 2 4 6
MV2D + Offline fine-tune 43.4 42.7 41.0 39.9
TTC-MV2D 51.0 50.7 50.2 49.6
In this section, we compare our TTC method with the approach that collects missing objects during inference and subsequently fine-tunes with the collected test-time 2D ground truth. This is another approach for utilizing test-time human feedback, though with delays in further model fine-tuning.
We select MV2D [57] as the baseline since it relies on 2D detection results for 3D object detection, thus allowing it to utilize 2D human feedback annotations to fine-tune. Additionally, we collect 2D feedback from various frame intervals, as users cannot provide feedback at every frame. For MV2D, we use all the 2D box annotations collected at specified intervals for fine-tuning. For the TTC-MV2D, we update the prompt buffer at these specified intervals.
As shown in Table 11, TTC-MV2D outperforms the MV2D model fine-tuned with 2D feedback. Notably, the performance of MV2D fine-tuned with human feedback from larger frame intervals, which means less frequent human feedback, declines significantly compared to models fine-tuned with feedback from every frame. In contrast, the TTC system allows the MV2D to perform effectively even with less frequent prompts of missed detections. This finding highlights the practicality of the TTC design in utilizing a single corresponding visual prompt for streaming data, as it is impractical for users to provide 2D prompts at each frame.
Appendix E Qualitative Results
We provide extensive visualizations to demonstrate the versatility of the TTC system across diverse scenarios:
  • In Section E.1 and Section E.2, we fix the prompt buffer with visual prompts of either labeled or novel, unlabeled 3D objects from the nuScenes dataset to detect targets in the nuScenes images.
  • In Section E.3, we test the performance with prompt buffer containing visual prompts in styles differing from the training distribution, such as Lego.
  • In Section E.4, we visualize the similarity maps on out-of-domain images, including YouTube driving videos and Internet-sourced visual prompts, demonstrating the generalization of the visual prompt alignment and the effectiveness of our TTC in reducing driving risks in non-standard scenarios.
For all visualizations, the fixed prompt buffer is shown in the first row. These comprehensive evaluations highlight the versatility of the TTC system in leveraging diverse visual prompts for 3D detection. All results are conducted with TTC-MonoDETR.
E.1 In-domain Visual Prompts on nuScenes “Seen” Objects
This visualization focuses on the in-domain detection performance of the TTC system on the nuScenes dataset. We utilize visual prompts from labeled objects in the nuScenes dataset, and demonstrate the system’s ability to effectively detect and track these target objects across different frames, as shown in Figure 13 and Figure 14. The results illustrate that our TTC can accurately localize and consistently track the target objects of interest within the nuScenes scenarios, showcasing its effectiveness in handling in-domain visual prompts.
E.2 In-domain Visual Prompts on nuScenes “Unseen” Objects
This visualization focuses on the TTC’s ability to handle visual prompts of objects not labeled in the nuScenes dataset. As shown in Figure 15, Figure 16, and Figure 17, our method demonstrates its potential to detect and track novel, out-of-distribution objects with these unseen visual prompts.
The results illustrate the TTC system’s capability to go beyond the training distribution and effectively localize and track objects that were not part of the original labeled dataset. This showcases the versatility and generalization ability of the visual prompt-based framework, enabling the detection of previously unseen objects. These findings highlight the potential of the TTC system to continuously expand its object detection capabilities by incorporating user-provided visual prompts, even for objects that were not included in the initial training data.
E.3 Out-domain Visual Prompts on nuScenes “Seen” Objects
This visualization focuses on the TTC’s performance with visual prompts in styles different from the training distribution. As shown in Figure 18 and Figure 19, the model can effectively detect target objects using visual prompts in various views and styles that diverge from the original training data.
These results demonstrate the potential of the TTC system to be extended to customized online 3D detection scenarios, where users can provide arbitrary visual prompts to guide the detection of objects of interest. The model’s ability to handle prompts across diverse style domains highlights its flexibility and versatility, a key advantage for enabling user-centric, interactive 3D perception systems.
E.4 Out-domain Visual Prompts on Real-world Examples
This visualization examines the generalization and robustness of TTC detectors in aligning visual prompts with the corresponding target objects in the input video stream. We select challenging driving scenarios involving unexpected animals running into the path of the ego vehicle. This is aimed at demonstrating the TTC system’s capability in reducing online driving risks in such non-standard situations.
As shown in Figure 20 and Figure 21, our method can effectively localize non-expected animals with higher responses in the regions where animals located, even though it was trained solely on the nuScenes dataset. Figure 22, Figure 23, Figure 24, Figure 25 further present examples of detection in a real-world scenario containing both vehicles and animals. Our method can detect all objects with their visual prompts simultaneously.
This further exemplifies the strong generalization capability of our TTC system, underscoring its potential for effectively handling challenging, edge-case scenarios on roads. The ability to accurately detect and localize unexpected objects beyond the training distribution highlights the robustness of the proposed approach, a key requirement for reliable autonomous driving systems.
Appendix F License of Assets
The adopted nuScenes dataset [1] is distributed under a CC BY-NC-SA 4.0 license. We implement the model based on mmDet3D codebase [8], which is released under the Apache 2.0 license.
We will publicly share our code and models upon acceptance under Apache License 2.0.
Figure 13: Visualizations on nuScenes scenarios with in-domain visual prompts of labeled objects. TTC system enables continuous 3D detection and tracking based on visual prompts. The images in the first row indicate the visual prompts in prompt buffer, and images in other rows represent 3D detection results prompted by the corresponding visual prompts. Different identities are indicated with different colors.
Figure 14: Visualizations on nuScenes scenarios with in-domain visual prompts of labeled objects. TTC system enables continuous 3D detection and tracking based on visual prompts. The images in the first row indicate the visual prompts in prompt buffer, and images in other rows represent 3D detection results prompted by the corresponding visual prompts. Different identities are indicated with different colors.
Figure 15: Visualizations on nuScenes scenarios with in-domain visual prompts of un-labeled objects. TTC system enables 3D detection and tracking of “novel” objects unseen during training. The image in the first row indicates the visual prompt in the prompt buffer, and images in other rows represent 3D detection results prompted by the corresponding visual prompts. Interestingly, with the one-to-N mapping mechanism of the visual prompt alignment, TTC system can detect multiple objects with similar visual descriptions to the visual prompt simultaneously.
Figure 16: Visualizations on nuScenes scenarios with in-domain visual prompts of un-labeled objects. TTC system enables 3D detection and tracking of “novel” objects unseen during training. The image in the first row indicates the visual prompt in the prompt buffer, and images in other rows represent 3D detection results prompted by the corresponding visual prompts.
Figure 17: Visualizations on nuScenes scenarios with in-domain visual prompts of un-labeled objects. TTC system enables 3D detection and tracking of “novel” objects unseen during training. The image in the first row indicates the visual prompt in the prompt buffer, and images in other rows represent 3D detection results prompted by the corresponding visual prompts. Interestingly, with the one-to-N mapping mechanism of the visual prompt alignment, TTC system can detect multiple objects with similar visual descriptions to the visual prompt simultaneously.
Figure 18: Visualizations on nuScenes scenarios with out-domain visual prompts of labeled objects. TTC system can perform 3D detection and tracking via visual prompts with arbitrary styles (imagery style). The image in the first row indicates the visual prompt in the prompt buffer, and images in other rows represent 3D detection results prompted by the corresponding visual prompts. With the one-to-N mapping mechanism of the visual prompt alignment, TTC detectors can detect multiple objects with similar visual descriptions to the visual prompt at the same time.
Figure 19: Visualizations on nuScenes scenarios with out-domain visual prompts of labeled objects. TTC system can perform 3D detection and tracking via visual prompts with arbitrary styles (Lego style). The image in the first row indicates the visual prompt in the prompt buffer, and images in other rows represent 3D detection results prompted by the corresponding visual prompts. With the one-to-N mapping mechanism of the visual prompt alignment, TTC detectors can detect multiple objects with similar visual descriptions to the visual prompt at the same time.
Figure 20: Visualizations on real world scenarios with out-of-domain visual prompts. At here, we show the similarity map from visual prompt alignment with visual prompts of arbitrary objects unseen during training. Brighter colors highlight higher responses. As illustrated, our method works well in novel scenarios with visual prompts of arbitrary object-of-interests. Best viewed in color.
Figure 21: Visualizations on real world scenarios with out-of-domain visual prompts. At here, we show the similarity map from visual prompt alignment with visual prompts of arbitrary objects unseen during training. Brighter colors highlight higher responses. As illustrated, our method works well in novel scenarios with visual prompts of arbitrary object-of-interests. Best viewed in color.
Figure 22: Visualizations on real world scenarios with out-of-domain visual prompts. This figure demonstrates the TTC’s effectiveness in real-world 3D detection, even with visual prompts of unseen objects. demonstrate a case of real-world detection As illustrated, our method works well in real-world scenarios with visual prompts of unseen objects. Different object identities are indicated by distinct colors.
Figure 23: Visualizations on real world scenarios with out-of-domain visual prompts. This figure demonstrates a case of real-world 3D object detection. We provide the visual prompt from the same scene as the target object, though it is unseen during training, and our TTC system then detects the target objects in the subsequent video frames.
Figure 24: Visualizations on real world scenarios with out-of-domain visual prompts. This figure demonstrates a case of real-world 3D object detection. We provide a visual prompt from a separate image as the pre-defined prompt and visualize it at the beginning of the sequence. As described, our TTC system can successfully detect the ”Deer” object using this stylized deer prompt sourced from the internet. This example highlights the TTC’s capability to effectively leverage diverse, user-supplied visual prompts to accurately identify target objects, even when the prompts are not directly from the same scene.
Figure 25: Visualizations on real world scenarios with out-of-domain visual prompts. Another case demonstrates the use of an internet-sourced ”Deer” visual prompt to detect the deer in a different real-world scenario. As shown, our TTC system effectively detects the deer even when it is partially obscured by snow. This example further illustrates the robust performance of the TTC framework in accurately localizing target objects, even in challenging environmental conditions, by leveraging flexible visual prompts provided by users.

### v3

@@
-Test-time Correction with Human Feedback: An Online 3D Detection System via Visual Prompting
+Test-time Correction: An Online 3D Detection System via Visual Prompting
 Abstract
-This paper introduces Test-time Correction (TTC) system, a novel online 3D detection system designated for online correction of test-time errors via human feedback, to guarantee the safety of deployed autonomous driving systems.
-Unlike well-studied offline 3D detectors frozen at inference, TTC explores the capability of instant online error rectification.
-By leveraging user feedback with interactive prompts at a frame, e.g., a simple click or draw of boxes, TTC could immediately update the corresponding detection results for future streaming inputs, even though the model is deployed with fixed parameters.
-This enables autonomous driving systems to adapt to new scenarios immediately and decrease deployment risks reliably without additional expensive training.
-To achieve such TTC system, we equip existing 3D detectors with Online Adapter (OA) module, a prompt-driven query generator for online correction.
-At the core of OA module are visual prompts, images of missed object-of-interest for guiding the corresponding detection and subsequent tracking.
-Those visual prompts, belonging to missed objects through online inference, are maintained by the visual prompt buffer for continuous error correction in subsequent frames.
-By doing so, TTC consistently detects online missed objects and immediately lowers driving risks.
-It achieves reliable, versatile, and adaptive driving autonomy.
-Extensive experiments demonstrate significant gain on instant error rectification over pre-trained 3D detectors, even in challenging scenarios with limited labels, zero-shot detection, and adverse conditions.
-We hope this work would inspire the community to investigate online rectification systems for autonomous driving post-deployment.
-Code would be publicly shared.
-Figure 1: TTC system could instantly respond to missed objects via human feedback, detecting and tracking these targets in subsequent frames consistently.
-As depicted on the left, human feedback might come from any views of target objects, including missed objects in the current and/or previous frames, as well as diverse viewpoints across scenes, styles, sources, and poses.
-Such a system improves offline-trained 3D detectors by rectifying online driving behavior immediately, reducing safety risks via test-time correction.
-TTC enables a reliable and adaptive online autonomous driving system.
-Equal contribution.
+This paper introduces Test-time Correction (TTC), an online 3D detection system designed to rectify test-time errors using various auxiliary feedback, aiming to enhance the safety of deployed autonomous driving systems.
+Unlike conventional offline 3D detectors that remain fixed during inference, TTC enables immediate online error correction without retraining, allowing autonomous vehicles to adapt to new scenarios and reduce deployment risks.
+To achieve this, we equip existing 3D detectors with an Online Adapter (OA) module—a prompt-driven query generator for real-time correction.
+At the core of OA module are visual prompts: image-based descriptions of objects of interest derived from auxiliary feedback such as mismatches with 2D detections, road descriptions, or user clicks.
+These visual prompts, collected from risky objects during inference, are maintained in a visual prompt buffer to enable continuous correction in future frames.
+By leveraging this mechanism, TTC consistently detects risky objects, achieving reliable, adaptive, and versatile driving autonomy.
+Extensive experiments show that TTC significantly improves instant error rectification over frozen 3D detectors, even under limited labels, zero-shot settings, and adverse conditions.
+We hope this work inspires future research on post-deployment online rectification systems for autonomous driving.
+Index Terms:
+3D Object Detection, Test-time correction, Online detection system.
 1 Introduction
+Visual-based 3D object detection, which localizes and classifies 3D objects from visual imagery, plays a crucial role in autonomous driving systems.
+Visual autonomous driving frameworks [hu2023_uniad, hu2022stp3, casas2021mp3, Cui_2021] rely heavily on accurate 3D detection outcomes to predict future driving behaviors and plan the trajectory of the ego vehicle.
+Existing 3D object detectors [li2022bevformer, yang2023bevformerV2, liu2022petr, CaDDN, huang2021bevdet] typically follow an offline training and deployment pipeline.
+Once the model is trained and deployed on self-driving cars, it is expensive to update new behaviors, e.g., another turn of offline re-training or fine-tuning.
+That is, when the system fails to perceive important objects or fails in novel scenarios due to a domain shift, these offline solutions cannot update themselves online to rectify mistakes immediately and detect missed objects.
+Such a caveat poses significant safety risks to reliable driving systems, e.g., dangerous driving behaviors such as improper lane changing, turning, or even collisions.
 Figure 1: Comparison of Error Correction between the conventional offline loop (left) and the new proposed online TTC System (right).
@@
 While TTC system additionally enables deployed 3D detectors with on-the-fly error rectification ability.
-Visual-based 3D object detection, which localizes and classifies 3D objects from visual imagery, plays a crucial role in autonomous driving systems.
-Visual autonomous driving frameworks [20, 19, 3, 9] rely heavily on accurate 3D detection outcomes to predict future driving behaviors and plan the trajectory of the ego vehicle.
-Existing 3D object detectors [27, 62, 35, 46, 21] typically follow an offline training and deployment pipeline.
-Once the model is trained and deployed on self-driving cars, it is expensive to update new behaviors, e.g., another turn of offline re-training or fine-tuning.
-That is, when the system fails to perceive important objects or fails in novel scenarios due to a domain shift, these offline solutions cannot update themselves online to rectify mistakes immediately and detect missed objects.
-Such a caveat poses significant safety risks to reliable driving systems, e.g., dangerous driving behaviors such as improper lane changing, turning, or even collisions.
 To guarantee safety, we argue that 3D detectors deployed on autonomous driving systems can rectify missed detection on the fly during test time.
-As depicted in Figure 2 (left), for error rectification, existing 3D detectors rely on offline pipelines, encompassing a full-suite procedure of data collection, annotation, training, and deployment.
+As depicted in Figure˜1 (left), for error rectification, existing 3D detectors rely on offline pipelines, encompassing a full-suite procedure of data collection, annotation, training, and deployment.
 This requires significant human workloads and resources for labeling and re-training, days or even weeks to fulfill.
 Other than the offline pipeline to improve models, we desire deployed 3D detectors also capable of test-time correction since such delays in offline updating are unacceptable when facing risks on the road, where safety is of the utmost priority.
-In this work, we explore a new 3D detection system capable of Test-Time error Correction based on human feedback online, namely TTC, akin to how human drivers respond, as shown in Figure 2 (right).
-The system is designed to address missed detection issues in autonomous driving scenarios, enabling existing 3D detectors to instantly correct errors through online human input, thereby preventing the risks of erroneous decisions caused by missed detections.
-Inspired by the principles of In-context Learning Customization [58, 43] in large language models (LLMs) [40, 52], we achieve the TTC system by leveraging images of missed objects collected from human feedback as context prompts.
-These prompts assist deployed 3D detectors in identifying and localizing previously unrecognized objects in later streaming input frames, without extra parameter updates during testing.
-The proposed TTC includes two components: Online Adapter (OA) that enables 3D detectors with visual promotable ability, and a visual prompt buffer that records missing objects.
-The core design is “visual prompts”, the visual object representation derived from human feedback.
-Existing promptable 3D detection methods typically utilize text, boxes, or clicks as prompts [4, 7, 66, 15, 2].
-However, text prompts can be ambiguous and may not describe the target objects effectively.
-Meanwhile, box and point prompts struggle to handle streaming data.
+In this work, we propose a novel 3D detection system, TTC, that provides online Test-Time Correction by leveraging various auxiliary information (e.g., 2D detections, road descriptions, or human feedback), as shown in Figure˜1 (right).
+The system is designed to address missed detection issues in autonomous driving scenarios, enabling existing 3D detectors to leverage various real-time feedback from multiple sources for instant error correction, thereby reducing the likelihood of erroneous decisions caused by undetected objects.
+Inspired by the principles of In-context Learning Customization [cot, llm2023peng] in large language models (LLMs) [gpt4, gemini], we implement test-time correction by converting auxiliary information into prompts and feeding them into the system for online adaptation.
+The proposed TTC includes two components: Online Adapter (OA) that enables 3D detectors with visual promotable ability, and a visual prompt buffer that records risky objects.
+The core design is “visual prompts”, the visual object representation derived from various auxiliary sources.
+Existing promptable 3D detection methods typically utilize text, boxes, or clicks as prompts [chen2024ll3da, choi2024idet3d, zhang2023sam3d, guo2024sam2point, cao2024coda].
+However, text prompts often require integration with large language models, and supplying a text prompt on a per-frame basis results in significant latency.
+Meanwhile, box and point prompts on specific frames struggle to handle streaming data.
 These limitations indicate that such prompts are inadequate for real-time autonomous driving tasks.
-Visual prompts cover arbitrary imagery views of target objects, i.e., views in different zones, styles, and timestamps (Figure 3), indicating the identity of target objects.
+Visual prompts cover arbitrary imagery views of target objects, i.e., views in different zones, styles, and timestamps (Figure˜2), indicating the identity of target objects.
+Once generated, visual prompts can operate on subsequent frames, accurately describing and tracking target risky objects.
 With visual prompts, OA module generates corresponding queries, locates corresponding objects within streaming inputs, and facilitates 3D detectors to output 3D boxes.
-To enable consecutive error rectification for video streaming, we design a dynamic visual prompt buffer to maintain visual prompts of all past unrecognized objects.
-In each iteration, we use all visual prompts in this buffer as inputs of the TTC system.
-This enables the continuous detection and tracking of all previously missed objects, redressing online errors for streaming input effectively.
-Further, to refrain the buffer from undesirable expansion, we introduce a “dequeue” operation to ensure its bounded size, allowing for consecutive rectification without excessive overhead.
+To enable consecutive error rectification for video streaming, we design a dynamic visual prompt buffer to maintain visual prompts of all risky objects.
+In each iteration, we use all visual prompts in this buffer as inputs to the TTC system.
+This enables the continuous detection and tracking of all risky objects, redressing online errors for streaming input effectively.
+Further, to prevent the buffer from undesirable expansion, we introduce a “dequeue” operation to ensure its bounded size, allowing for consecutive rectification without excessive overhead.
 Figure 2: Visual prompts could be arbitrary views of objects, across zones, styles, timestamps, etc.
-We conduct experiments on the nuScenes dataset [1].
+We conduct experiments on the nuScenes dataset [nuscenes2019].
 Given the novel setting of our TTC system, the assessment focuses on its abilities for instant online error correction.
@@
 Remarkably, the TTC system significantly improves the offline 3D detectors for test-time performance.
-Specifically, with test-time rectification, TTC improves offline monocular [67], multi-view [57], and BEV detectors [62] by 5.0%, 12.7%, and 12.1% EDS  [^2] EDS is a class-agnostic version of nuScenes Detection Score (NDS) [1], treating all objects as class-agnostic entities and ignoring the velocity and attribute to evaluate out-of-distribution 3D detection.
-, respectively.
-Second, on challenging scenarios, the TTC system exhibits even more substantial gains, with improvements of 14.4%, 21.6%, 13.6%, and 4.7% EDS on tasks like distant 3D detection and vehicle-focused detection with limited annotations, zero-shot extensions, as well as scenarios with domain shifts, respectively.
+Specifically, with test-time rectification, TTC improves offline monocular [zhang2022monodetr], multi-view [wang2023object], and BEV detectors [yang2023bevformerV2] by 2.4%, 9.7%, and 9.3% mAP, respectively.
+Second, on challenging scenarios, the TTC system exhibits even more substantial gains, with improvements of 10.1%, 20.0%, 7.2%, and 3.8% mAP on tasks like distant 3D detection and vehicle-focused detection with limited annotations, zero-shot extensions, as well as scenarios with domain shifts, respectively.
 This comprehensive evaluation highlights the versatility and adaptability of TTC, which can effectively rectify online errors and maintain robust 3D detection even under limited data, category shifts, and environmental changes.
@@
 In this paper, we introduce the task of online 3D detection, which focuses on the instant rectification of errors during the online testing phase of 3D detectors.
-The primary goal is to enable the continuous detection of objects missed by offline-trained 3D detectors, without requiring additional training.
-This work is closely related to several areas, including 3D detection, tracking, test-time adaption and interactive vision models.
-Here, we discuss 3D detection and tracking, while leave others in the supplementary material Appendix B.
-3D Object Detection.
+The primary goal is to enable the continuous detection of objects missed by offline-trained 3D detectors, without requiring additional training after deployment.
+This work is closely related to several areas, including 3D detection, tracking, test-time adaptation, and interactive vision models.
+2.1 3D Object Detection
 3D object detection is a cornerstone task in autonomous driving.
-Traditional 3D detectors primarily focus on the performance of pre-trained models [50, 25, 65, 44, 27].
+Traditional 3D detectors primarily focus on the performance of pre-trained models [PVRCNN, lang2018pointpillars, yin2021center, philion2020lift, li2022bevformer].
 Once deployed, the model’s output is fixed during inference, with no ability to intervene in real time to correct missed detections.
-In recent years, some works have explored instruction-based detection systems, offering interfaces that accept human prompts [4, 7, 66, 15, 70, 2, 55].
+In recent years, some works have explored instruction-based detection systems, offering interfaces that accept prompts [chen2024ll3da, choi2024idet3d, zhang2023sam3d, guo2024sam2point, zhu2025unifying, cao2024coda, wang2024omnidrive].
 However, these systems typically rely on boxes, points, or language as interfaces.
-Boxes and points struggle to handle temporal data [7, 66, 15], while language prompts often depend on LLMs [4, 70, 2, 55], leading to higher latency and potential ambiguities.
+Boxes and points struggle to handle temporal data [choi2024idet3d, zhang2023sam3d, guo2024sam2point], while language prompts often depend on LLMs [chen2024ll3da, zhu2025unifying, cao2024coda, wang2024omnidrive], leading to higher latency and potential ambiguities.
 In contrast, online 3D detection focuses on real-time error correction during sequential detection.
 By utilizing visual prompts, we enable accurate correction of target objects in streaming data instantly.
-Target Object Tracking.
-Tracking is another direction closely related to online 3D detection [28, 41, 42, 68, 18].
-Since our approach uses visual prompts to identify targets, it closely resembles single object tracking (SOT) in terms of its objectives [61, 5, 10, 11].
+2.2 Target Object Tracking
+Tracking is another direction closely related to online 3D detection [pnpnet, pang2021simpletrack, Multiple3Dtracking, zhang2022mutr3d, Hu2021QD3DT].
+Since our approach uses visual prompts to identify targets, it closely resembles single object tracking (SOT) in terms of its objectives [yan2021learning, chen2023seqtrack, cui2022mixformer, cui2024mixformerv2].
 While current SOT methods based on images are limited to generating 2D detection outputs and presume the object query as the initial frame of the ongoing video, our TTC system stands out.
 It can directly identify and track the associated 3D bounding boxes using visual prompts, even when these prompts originate from varied sources, scenes, and zones with distinct styles.
-3 Test-time Correction with Human Feedback
+2.3 Interactive Vision Models
+Interactive vision models are designed for tasks based on user inputs.
+As one of the fundamental tasks in computer vision, they are extensively researched with numerous breakthroughs [lazysnap, focalclick, simpleclick, deepinteract, randomwalk].
+Particularly, the advent of the Segment Anything Model (SAM) [kirillov2023segany] has sparked a surge of progress, with applications spanning reconstruction [shen2023anything3d], detection [ren2024grounded, yang2023track], segmentation [zou2023segment], image editing [gao2023editanything], and more.
+Compared to existing models, which typically rely on prompts such as clicks, boxes, or scribbles, in this work, we study visual prompts, a new prompt referring to the actual images of objects with arbitrary poses and styles.
+Visual prompts enable continuous detection and tracking of target objects in streaming videos, facilitating the immediate correction of failed detections at test time.
+This represents a departure from traditional prompt types, offering more natural and dynamic interactions with the visual content.
+2.4 Test-time Adaptation
+Test-time adaptation aims to improve model performance on test data by adapting the model to test samples, even in the presence of data shifts [zhang2023adanpc, niu2022efficient, niutowards, schneider2020improving, wang2020tent, liu2021ttt++].
+Most TTA methods update network parameters during testing—via entropy minimization, self-supervised objectives, or batch-norm statistic adaptation.
+More recently, several works have brought TTA to object detection [ruan2024fully, vs2023towards, cao2024exploring], and to 3D detection in particular (e.g., MonoTTA [lin2025monotta], Reg-TTA3D [yuan2024reg], and MOS [chenmos]).
+These approaches typically fine-tune on streaming target batches across domains or perform input-conditioned model selection/mixture, still relying on weight updates or model switching at test time.
+In contrast, our method does not modify the original model parameters.
+Instead, it performs parameter-free, prediction-level rectification driven by various auxiliary feedback, enabling instance-level online correction without retraining or compromising source-domain accuracy.
+3 Test-time Correction
 Figure 3: Overall Framework.
@@
 It supports 3D detection from any combination of four prompts, i.e., object $\mathcal{P}_{o}$, box $\mathcal{P}_{b}$, point $\mathcal{P}_{p}$, and novel visual prompts $\mathcal{P}_{v}$, arbitrary views of target objects across scenarios and timestamps.
-In this section, we elaborate on our TTC, an online test-time error rectification system for 3D detection to detect and track missed objects during on-road inference with the guidance of human feedback.
-We start with an overview in Section 3.1, then delve into OA module and the visual prompt buffer in Section 3.2, and Section 3.3, respectively.
+In this section, we elaborate on our TTC, an online test-time error rectification system for 3D detection to detect and track missed objects during on-road inference with the guidance of various feedback.
+We start with an overview in Section˜3.1, then delve into OA module and the visual prompt buffer in Section˜3.2, and Section˜3.3, respectively.
 3.1 Overview
-We convert online human feedback, i.e., clicks, boxes, or images from the Internet, into a uniform representation called “visual prompts”, image descriptions of target objects.
+We convert various online auxiliary feedback, i.e., such as human-provided clicks and boxes, road descriptions, and misaligned detection results compared with 2D detectors, into a uniform representation called “visual prompts”, image descriptions of missed or risky objects.
 Such image-based descriptions can cover arbitrary views of objects, including pictures taken from diverse zones, weather, timestamps, or even from out-of-domain sources such as stylized Internet images.
-Upon visual prompts, the TTC system, a recurrent framework, is designed to engage in sustained interaction with human users, continuously learning to detect and track new objects.
-As Figure 4 (left) shows, it comprises two key components: 1) TTC-3D Detector, any 3D detector equipped with OA module for in-context 3D detection and tracking via visual prompts, and 2) an extendable visual prompt buffer storing visual prompts of all previously missed objects, enabling continuous online error rectification.
-During online inference after being deployed on cars, whenever an error occurs, i.e., miss an object, users can add the unrecognized object to visual prompt buffer by clicking on it in the image.
-The model then detects the corresponding 2D boxes based on the user-provided clicking prompt and updates the visual prompt buffer with the associated image patch.
-In subsequent frames, TTC-3D Detector leverages the stored visual prompts to detect and track previously missed 3D objects.
+Based on visual prompts, the TTC system is designed as a real-time framework capable of continuously detecting and tracking risky objects, even when the raw feedback may suffer from high latency.
+As Figure˜3 (left) shows, it comprises two key components: (i) TTC-3D Detector, any 3D detector equipped with OA module for in-context 3D detection and tracking via visual prompts, and (ii) an extendable visual prompt buffer storing visual prompts of all previously missed objects, enabling continuous online error rectification.
+During online inference after deployment in vehicles, the model can leverage auxiliary information by converting it into visual prompts and adding them to the visual prompt buffer.
+For user clicks, the model first detects the corresponding 2D boxes based on the user-provided point prompt and updates the visual prompt buffer with the associated image patch.
+For user-drawn boxes or cases involving misaligned 2D detection results, the corresponding image regions can be directly pre-processed via cropping.
+For road descriptions, relevant visual prompts can be retrieved via online search based on the provided textual descriptions.
+TTC-3D Detector leverages the stored visual prompts to detect and track risky 3D objects.
 This enables instant error correction and continuous improvement of 3D detection during online operation.
@@
 OA module is designed as a plug-and-play module.
-It receives human prompts and transforms them into queries that can be processed by traditional detectors.
-When trained with OA module, traditional detectors can understand human prompts to produce detection results, enabling online error correction without the need for parameter updates.
+It receives prompts and transforms them into queries that can be processed by traditional detectors.
+When trained with OA module, traditional detectors can understand various prompts to produce detection results, enabling online error correction without the need for parameter updates.
 Prompt-driven Query Generating.
-As shown in Figure 4 (right), OA module is flexible to handle four prompts in different forms: object query prompts for traditional offline 3D detection, box and point prompts for collecting test-time feedback, and visual prompts for consistent error correction.
+As shown in Figure˜3 (right), OA module is flexible to handle four prompts in different forms: object query prompts for traditional offline 3D detection, box and point prompts for feedback converting, and visual prompts for consistent error correction.
 Specifically, these prompts are processed as follows:
-  • For object prompts $\mathcal{P}_{o}$, the OA module generates a set of learnable embeddings as queries, akin to traditional 3D detectors, which are updated during training as demonstrated in previous works [27, 56, 67].
-  • For box $\mathcal{P}_{b}$ and point $\mathcal{P}_{p}$ prompts, the OA module encodes them with their location and shape, representing them as Fourier features [51].
-  • For visual prompts $\mathcal{P}_{v}$, the OA module first extracts the visual prompt features $\mathcal{Z}_{v}$ by an encoder [17], then localizes their corresponding objects within the input images $\mathcal{X}_{v}$, and finally adds their features with the Fourier positional encoding as subsequent inputs:
+  • For object prompts $\mathcal{P}_{o}$, the OA module generates a set of learnable embeddings as queries, akin to traditional 3D detectors, which are updated during training as demonstrated in previous works [li2022bevformer, detr3d, zhang2022monodetr].
+  • For box $\mathcal{P}_{b}$ and point $\mathcal{P}_{p}$ prompts, the OA module encodes them with their location and shape, representing them as Fourier features [tancik2020fourfeat].
+  • For visual prompts $\mathcal{P}_{v}$, the OA module first extracts the visual prompt features $\mathcal{Z}_{v}$ by an encoder [ResNet], then localizes their corresponding objects within the input images $\mathcal{X}_{v}$, and finally adds their features with the Fourier positional encoding as subsequent inputs:
     $\mathcal{P}_{v}=\texttt{FourierPE}(\texttt{Align}(\mathcal{Z}_{v},\mathcal{Z}))+\mathcal{Z}_{v},$ (1)
@@
 We perform the Align operation to localize target objects in the input images, by visual prompts.
-To handle flexible visual prompts, which can be image descriptions of objects in any view, scene, style, or timestamp, we employ contrastive mechanisms [59, 16] as the key design to retrieve target objects at different styles.
+To handle flexible visual prompts, which can be image descriptions of objects in any view, scene, style, or timestamp, we employ contrastive mechanisms [wu2018unsupervised, he2020moco] as the key design to retrieve target objects at different styles.
 This allows the module to detect the target objects effectively, even when they exhibit diverse visual styles and appearances.
-Specifically, we use two multi-layer perceptrons (MLPs) to first align the channels of image features $\mathcal{Z}$ and visual prompt features $\mathcal{Z}_{v}$.
-We then compute the dot product between the aligned feature maps to obtain a similarity map.
-To further retrieve the coordinates of target objects from the similarity map, we multiply it with the original image features $\mathcal{Z}$, and apply another MLP with channels of 2 to regress the spatial positions $\mathcal{X}_{v}$ of target objects.
+Figure 4: Concrete mechanism of visual prompt alignment.
+This figure illustrates monocular input.
+When multi-view images are employed, this alignment operation flattens the different views and still generates NN peak candidate positions.
+As described in Figure˜4, given the visual prompt $\mathcal{P}_{v}$, we first extract the visual prompt features $\mathcal{Z}_{v}$ using a lightweight encoder, e.g., ResNet18 [ResNet] in our implementation.
+Then, we use two separate 2-layer perceptrons, each with 128 and 64 output channels, to align the channel dimensions of the prompt features and the input image features.
+This results in the aligned feature maps $\hat{\mathcal{Z}_{v}}\in\mathbb{R}^{M\times 64}$ and $\hat{\mathcal{Z}}\in\mathbb{R}^{KHW\times 64}$, where $M$ is the number of visual prompts, $K$ is the number of image views, $H$ and $W$ are the spatial dimensions of image features.
+Finally, we compute the cosine similarity ($\odot$ in Figure˜4) between $\hat{\mathcal{Z}}_{v}$ and $\hat{\mathcal{Z}}$ to obtain the similarity map $\mathcal{S}\in\mathbb{R}^{M\times KHW}$, which encodes the alignments between the visual prompts and image features.
 Instance Ambiguity & Loss.
 Sometimes, visual prompts might exhibit instance ambiguity, where multiple objects in the image match the visual descriptions of prompts.
-For example, suppose the visual prompt describes a traffic cone and several similar-looking traffic cones present in the input images.
-It can be challenging to uniquely identify the specific object-of-interest.
+For example, suppose the visual prompt describes a traffic cone and several similar-looking traffic cones present in the input images (highlighted by “orange box” in Figure˜4).
+It can be challenging to uniquely identify the specific object of interest.
 For such cases, we design to retrieve all objects with similar identities to the visual prompt.
@@
 For similarity supervision, we generate binary segmentation labels based on the ground-truth 2D boxes.
-Focal loss [30] and Dice loss [37] are used for optimization.
-To supervise the visual prompt localization $\mathcal{X}_{v}$, we use the Smooth-$l$1 loss with the target as the center coordinates of ground-truth 2D bounding boxes.
-For dealing with instance ambiguity, we refer to SAM [24] and only backpropagate the sample with the minimum localization loss during each training iteration.
-For more details, please refer to the supplementary material Section C.2.
+Focal loss [FocalLoss] and Dice loss [Milletari2016vnet] are used for optimization.
+To supervise the visual prompt localization $\mathcal{X}_{v}$, we use the Smooth-ll1 loss with the target as the center coordinates of ground-truth 2D bounding boxes.
+For dealing with instance ambiguity, we refer to SAM [kirillov2023segany] and only backpropagate the sample with the minimum localization loss during each training iteration.
 Model Design.
-The overall mechanism of TTC-3D Detector is depicted in Figure 4 (right).
+The overall mechanism of TTC-3D Detector is depicted in Figure˜3 (right).
 Based on any traditional offline-trained 3D detector, BEV detector, or monocular detector, we integrate OA module and train it to be promptable.
 Specifically, OA module takes features extracted by the image encoder of the corresponding 3D detector, along with various forms of prompts as inputs.
-It encodes these prompts and generates a series of queries, represented as $\mathcal{P}=\{\mathcal{P}_{o};<\mathcal{P}_{b},\mathcal{P}_{p},\mathcal{P}_{v}>\}$, where <>expectation<...>< … > denotes an arbitrary combination of different prompts.
-These queries are then fed into the transformer decoder of the 3D detector to output 3D boxes following human feedback.
+It encodes these prompts and generates a series of queries, represented as $\mathcal{P}=\{\mathcal{P}_{o};<\mathcal{P}_{b},\mathcal{P}_{p},\mathcal{P}_{v}>\}$, where <><...> denotes an arbitrary combination of different prompts.
+These queries are then fed into the transformer decoder of the 3D detector to output 3D boxes.
 3.3 Visual Prompt Buffer
-Visual prompt buffer is a queue that stores user-provided visual prompts of missed objects during online inference.
-In the default setting, these visual prompts are the 2D output regions from the model in previous frames, based on the user’s point or box prompts at that time.
-The flexibility of the TTC system also allows users to freely select visual prompts, either image contents from the current scene or customized objects from the Internet.
+Visual prompt buffer is a queue that stores visual prompts of missed objects during online inference.
+In the default setting, these visual prompts are the 2D regions of risky objects in previous frames.
+The TTC system is also flexible enough to allow freely uploading visual prompts, which can be either image content from the current scene or customized images sourced from the Internet based on road descriptions or user-specified requests.
 This versatility makes the TTC system applicable to a wide spectrum of scenarios.
 Dequeue.
-To prevent the buffer from growing indefinitely, we design a “dequeue” mechanism.
-We filter out visual prompts with low confidence, which likely no longer appear in the scene.
-We also use the intersection-over-union (IoU) between predictions to identify and remove redundant visual prompts.
-This dynamic update and maintenance ensure a balance of TTC system between latency and accuracy by adaptively incorporating and pruning feedback online.
+To prevent unbounded growth, we adopt a dequeue policy that combines score-based eviction and redundancy pruning.
+During training, we mix in negative visual prompts—patches that do not correspond to any target in the current frame (e.g., background crops or off-timestamp instances)—and supervise the model to assign them low confidence.
+This calibration causes non-existent or contradictory prompts to be naturally down-weighted at test time, enabling confidence-based removal.
+Concretely, we first filter out prompts whose associated detections consistently fall below a confidence threshold.
+We then suppress near-duplicates by measuring intersection-over-union (IoU) between predictions and removing redundant prompts.
+This dynamic maintenance balances latency and accuracy by adaptively incorporating informative prompts while pruning stale or duplicate ones online.
 4 Experiments
-We now proceed to evaluate our TTC system from two key aspects:
+We proceed to evaluate our TTC system from two key aspects:
   • How does TTC perform with various adapted offline 3D detectors for instant online error rectification?
@@
 4.1 Experimental Setup
-Dataset. Experiments are done on nuScenes dataset [1] with 1,000 autonomous driving sequences, one of the most popular datasets for autonomous driving research.
+Dataset. Experiments are done on the nuScenes dataset [nuscenes2019] with 1,000 autonomous driving sequences, one of the most popular datasets for autonomous driving research.
 Tasks.
@@
 The first experiment aims to test the TTC system in boosting the performance of offline-trained 3D detectors via online error rectification.
-We conduct this verification by applying TTC system to various established offline 3D detectors [67, 27, 62, 32, 54, 31, 57].
+We conduct this verification by applying TTC system to various established offline 3D detectors [zhang2022monodetr, li2022bevformer, yang2023bevformerV2, liu2024ray, wang2023streampetr, sparse4dv2, wang2023object].
 Then, we validate the TTC system to correct detection errors in out-of-training-distribution scenarios.
 To conduct a thorough quantitative analysis of this aspect, we established four tasks under different settings: (a.) discarding 80% labels of distant objects farther than 30m during training, and test the error corrections for detecting distant objects; (b.) discarding 80% labels of instances labeled as vehicle, including “car”, “truck”, “C.V.”, “bus”, and “trailer”, and test the error correction in vehicles; (c.) discarding all annotations of class “truck” and “bus”, and test the zero-shot ability of TTC on those discarded classes; (d.) discarding all training data of the scenario of “Nighttime” and “Rainy”, and test the improvements of TTC on scenarios with domain gap.
-For these tasks, we base TTC system on the monocular algorithm, MonoDETR [67], and test it under each setting separately, as the monocular setting demonstrates the most general applicability.
-Based on TTC-MonoDETR, we also present qualitative results to show the potential of TTC to address corner cases in challenging real-world scenarios through human feedback during test time.
+For these tasks, we base TTC system on the monocular algorithm, MonoDETR [zhang2022monodetr], and test it under each setting separately, as the monocular setting demonstrates the most general applicability.
+Based on TTC-MonoDETR, we also present qualitative results to show the potential of TTC to address corner cases in challenging real-world scenarios through auxiliary feedback during test time.
 Through these experiments, we demonstrate TTC as an effective system to adapt offline 3D detectors to challenging scenarios.
@@
 As we focus on enabling existing 3D detectors to rectify missing detections, both in-distribution and out-of-distribution, we adopt a class-agnostic setting.
-We remove all class annotations of 3D objects during training, treating all objects as entities [24, 45].
-For evaluation, we use the Entity Detection Score (EDS), a class-agnostic version of the nuScenes Detection Score (NDS) that emphasizes recall.
-A detailed analysis of the rationale behind EDS can be found in Appendix Section C.3.
-Implementation Details.
-We implement our method based on MMDetection3D codebase [8], and conduct all experiments on a server with 8$\times$ A100 GPUs.
-In OA module, we use a ResNet18 [17] to extract visual prompt features.
-All visual prompts are resized to $224\times 224$ before being sent to OA module.
-For training, we use AdamW [23, 36] with a batch size of 16.
-We initialize the learning rate as 2e-4, adjusted by the cosine annealing policy.
-When training point and box prompts, we simulate user inputs with noise by adding perturbations to the ground truth.
-To ensure the visual prompts are robust across multiple scenes, timestamps, and styles, we choose visual prompts of target objects not only from the image patches of the current frame but also randomly from previous and future frames within a range of ±$\pm 5$.
-Flip operations are used as data augmentations.
-During testing, we mimic user intervention.
-We simulate the user’s online missing feedback by comparing the distance between ground truth and detected 3D boxes.
-Ground-truths without any detection within 2m are considered missed and added into the prompt buffer.
-We remove redundant predictions by non-max-suppression (NMS) with the IoU threshold of 0.5 and set the classification confidence threshold at 0.3.
-Table 1: Effect of TTC system on various 3D detectors. TTC system effectively improves the test-time performance of offline-trained detectors during online inference without any extra training.
-Method Backbone Type mAP (%) $\uparrow$ EDS (%) $\uparrow$
-MonoDETR [67] R101 Monocular 37.9 38.2
-TTC-MonoDETR 42.6 (+5.7) 43.2 (+5.0)
-MV2D [57] R50 Multiview 38.9 38.3
-TTC-MV2D 51.0 (+12.1) 51.0 (+12.7)
-Sparse4Dv2 [31] R50 Multiview + Temporal 38.8 37.6
-TTC-Sparse4Dv2 53.5 (+14.7) 52.1 (+14.5)
-BEVFormer [27] R101 BEV + Temporal 36.5 35.8
-TTC-BEVFormer 47.8 (+11.3) 47.2 (+11.4)
-BEVFormerV2-t8 [62] R50 BEV +Temporal 39.6 38.9
-TTC-BEVFormerV2-t8 51.6 (+12.0) 51.0 (+12.1)
-RayDN [32] R50 BEV + Temporal 39.6 38.7
-TTC-RayDN 51.7 (+12.1) 50.8 (+12.1)
-StreamPETR [54] V2-99 BEV + Temporal 39.7 39.1
-TTC-StreamPETR 52.3 (+12.6) 51.5 (+12.4)
+We remove all class annotations of 3D objects during training, treating all objects as entities [kirillov2023segany, qi2022openworld].
+For evaluation, we use the Entity Detection Score (EDS), a class-agnostic version of the nuScenes Detection Score (NDS).
+[…]
+TABLE 1: Effect of TTC on various 3D detectors.
+TTC improves test-time performance of offline-trained detectors during online inference without extra training.
+To compare directly with the baselines’ official scores, we set the confidence threshold to 0.01.
+For false-alarm analysis, we report FP/frame and FN/frame.
+Method Backbone Type mAP (%) $\uparrow$ EDS (%) $\uparrow$ FP/frame@0.01 $\downarrow$ FN/frame@0.01 $\downarrow$
+MonoDETR [zhang2022monodetr] R101 Monocular 42.6 50.1 49.9 1.2
+TTC-MonoDETR 45.0 (+2.4) 51.7 (+1.6) 49.9 0.9
+MV2D [wang2023object] R50 Multiview 43.0 52.8 32.2 2.0
+TTC-MV2D 52.7 (+9.7) 57.2 (+4.4) 32.2 0.8
+Sparse4Dv2 [sparse4dv2] R50 Multiview + Temporal 44.4 53.2 42.7 1.9
+TTC-Sparse4Dv2 54.2 (+9.8) 57.8 (+4.6) 42.7 0.7
+BEVFormer [li2022bevformer] R101 BEV + Temporal 41.1 50.2 34.2 1.9
+TTC-BEVFormer 50.2 (+9.1) 54.4 (+4.2) 34.4 0.8
+BEVFormerV2-t8 [yang2023bevformerV2] R50 BEV + Temporal 45.6 54.1 42.2 1.7
+TTC-BEVFormerV2-t8 54.9 (+9.3) 58.6 (+4.5) 42.3 0.6
+RayDN [liu2024ray] R50 BEV + Temporal 44.9 53.6 39.8 1.8
+TTC-RayDN 54.4 (+9.5) 58.1 (+4.5) 39.8 0.7
+StreamPETR [wang2023streampetr] V2-99 BEV + Temporal 45.0 53.8 42.3 1.8
+TTC-StreamPETR 51.6 (+6.6) 55.6 (+1.8) 42.4 0.7
 TABLE 2: Experiments on out-of-training-distribution scenarios. TTC system achieves substantial gains with limited or even no labels under different challenging test cases.
@@
 mAP (%) EDS (%) mAP (%) EDS (%)
-Point 41.6 40.8 17.8 19.4
-Box 44.3 43.7 19.2 21.7
-Visual 42.6 42.0 18.3 21.6
-MonoDETR 31.4 31.6 0.0 0.0
-TTC-MonoDETR 40.2 40.1 11.0 14.4
-$\Delta$ +8.8 +8.5 +11.0 +14.4
+Point 41.6 52.5 17.8 35.0
+Box 44.3 54.3 19.2 36.3
+Visual 42.6 53.0 18.3 36.1
+MonoDETR 31.4 47.6 0.0 0.0
+TTC-MonoDETR 39.6 52.1 10.1 27.9
+$\Delta$ +8.2 +4.5 +10.1 +27.9
 (b) Vehicle-focused rectification. Effect of TTC system on vehicle objects with 20% annotations.
@@
 mAP (%) EDS (%) mAP (%) EDS (%)
-Point 38.0 36.9 33.5 36.6
-Box 41.2 40.0 36.7 39.6
-Visual 39.2 38.4 34.6 38.5
-MonoDETR 17.6 16.1 2.4 7.3
-TTC-MonoDETR 29.0 27.2 23.2 28.9
-$\Delta$ +11.4 +11.1 +20.8 +21.6
+Point 38.0 48.2 33.5 52.8
+Box 41.2 50.3 36.7 54.4
+Visual 39.2 47.7 34.6 51.8
+MonoDETR 17.6 29.0 2.4 36.0
+TTC-MonoDETR 28.5 35.4 22.4 44.2
+$\Delta$ +10.9 +6.4 +20.0 +8.2
 (c) Novel object rectification. Effect of TTC system on objects of novel classes unseen in the training set.
@@
 mAP (%) EDS (%) mAP (%) EDS (%)
-Point 35.2 35.6 11.9 15.0
-Box 38.5 38.8 14.7 16.9
-Visual 35.4 36.0 11.2 15.6
-MonoDETR 28.3 29.6 0.0 0.0
-TTC-MonoDETR 34.8 36.1 8.5 13.6
-$\Delta$ +6.5 +6.5 +8.5 +13.6
+Point 35.2 48.7 11.9 33.3
+Box 38.5 50.9 17.0 36.6
+Visual 35.4 47.3 11.2 33.3
+MonoDETR 28.3 47.0 0.0 0.0
+TTC-MonoDETR 34.6 49.0 7.2 30.8
+$\Delta$ +6.3 +2.0 +7.2 +30.8
 (d) Domain shift rectification. Effect of TTC system on objects in scenarios with domain gap.
@@
 mAP (%) EDS (%) mAP (%) EDS (%)
-Point 39.0 38.2 30.5 30.7
-Box 42.6 41.5 33.4 33.6
-Visual 39.8 39.4 29.4 30.8
-MonoDETR 34.5 34.7 25.2 26.6
-TTC-MonoDETR 39.9 40.0 29.7 31.3
-$\Delta$ +5.4 +5.3 +4.5 +4.7
+Point 39.0 49.4 30.5 44.3
+Box 42.6 51.8 33.4 46.3
+Visual 39.8 48.6 29.4 43.0
+MonoDETR 34.5 48.4 25.2 43.1
+TTC-MonoDETR 39.5 51.8 29.0 44.5
+$\Delta$ +5.0 +3.4 +3.8 +1.4
 Figure 5: Qualitative visualization of real-world scenes (collected from YouTube).
@@
 Effectiveness of TTC in Test-time Error Correction.
-Test-time error correction ability without re-training is the core capability of TTC system.
-We verify this by incorporating traditional offline-trained 3D detectors into TTC system and compare the performance without test-time parameter updates.
-For thorough verification, we select various offline-trained 3D detectors, including monocular [67], multi-view [57, 31], and BEV ones [27, 62, 54, 32].
-As shown in Table 1, the TTC system substantially improves the test-time performance of offline-trained 3D detectors, e.g., 11.4% and 12.4% EDS improvements on BEVFormer and StreamPETR, without requiring additional training during inference.
+Test-time error correction ability without re-training is the core capability of the TTC system.
+We verify this by incorporating traditional offline-trained 3D detectors into the TTC system and compare the performance without test-time parameter updates.
+For thorough verification, we select various offline-trained 3D detectors, including monocular [zhang2022monodetr], multi-view [wang2023object, sparse4dv2], and BEV ones [li2022bevformer, yang2023bevformerV2, wang2023streampetr, liu2024ray].
+The online feedback is derived from the human clicks which is then converted into visual prompts.
+The feedback interval is set to 5 frames.
+As shown in Table˜1, the TTC system substantially improves the test-time performance of offline-trained 3D detectors, e.g., 9.1% and 6.6% mAP improvements on BEVFormer and StreamPETR, without requiring additional training after deployment.
 These demonstrate the effectiveness of the TTC system in instantly correcting test-time errors during online inference.
 Effectiveness of TTC in Out-of-Training Scenarios.
-Table 2 presents results to validate the TTC system in challenging and extreme scenarios.
+Table˜2 presents results to validate the TTC system in challenging and extreme scenarios.
 In these experiments, we base TTC system on MonoDETR for its simple monocular setting.
-“Point”, “Box”, and “Visual” in Table 2 are TTC-MonoDETR using point, box, or image patch of target objects in the input image as input prompts.
+“Point”, “Box”, and “Visual” in Table˜2 are TTC-MonoDETR directly using point, box, or image patch of target objects in the input image as input prompts, which are impractical unless automated annotation scenario.
 These are performance upper bounds as they receive human feedback in every frame.
-Table 2(a) evaluates the effectiveness of TTC system in rectifying detection errors on distant objects (beyond 30m).
+Compared with the baseline MonoDETR, “TTC-MonoDETR” in the table obtains visual prompts of missing objects from user at a feedback interval of 5 frames, enabling continuous correction without any test-time weight updates.
+Table˜II(a) evaluates the effectiveness of TTC system in rectifying detection errors on distant objects (beyond 30m).
 During training, 80% of the far-away annotations are removed, resulting in a 3D detector with poor long-range performance (0.0% mAP and EDS).
-Powered by the TTC system, the performance on these distant objects is instantly improved to 11.0% mAP and 14.4% EDS, without any extra training.
-The experiment is then extended to all vehicles in the nuScenes dataset, as shown in Table 2(b).
-The TTC achieves impressive performance gains of 20.8% mAP and 21.6% EDS.
-Furthermore, in Table 2(c) and Table 2(d), we evaluate the TTC system in two challenging scenarios: encountering unseen objects not present in the training data; handling domain shifts, such as transitioning from sunny to rainy or nighttime conditions.
-In Table 2(c), the TTC system shows effectiveness in successfully detecting novel objects not labeled in the training set, achieving 8.5% mAP and 13.6% EDS on these novel objects, significantly improving the offline-trained baseline with 0.0% mAP and EDS.
-In Table 2(d), we find TTC also works well for online error rectification when driving into scenarios with domain shifts, providing 4.5% mAP and 4.7% EDS improvements.
-Regarding qualitative results, Figure 5 further illustrates a case of the zero-shot capability in the real-world.
-Despite being trained solely on the nuScenes dataset, TTC-MonoDETR can detect a Deer using a visual prompt of another deer from a different viewpoint, which is extremely challenging for traditional offline detectors.
-More qualitative examples can be found in supplementary material Appendix E.
+Powered by the TTC system, the performance on these distant objects is instantly improved to 10.1% mAP and 27.9% EDS, without any extra training.
+The experiment is then extended to all vehicles in the nuScenes dataset, as shown in Table˜II(b).
+The TTC achieves impressive performance gains of 20.0% mAP and 8.2% EDS.
+Furthermore, in Table˜II(c) and Table˜II(d), we evaluate the TTC system in two challenging scenarios: encountering unseen objects not present in the training data; handling domain shifts, such as transitioning from sunny to rainy or nighttime conditions.
+In Table˜II(c), the TTC system shows effectiveness in successfully detecting novel objects not labeled in the training set, achieving 7.2% mAP and 30.8% EDS on these novel objects, significantly improving the offline-trained baseline with 0.0% mAP and EDS.
+In Table˜II(d), we find TTC also works well for online error rectification when driving into scenarios with domain shifts, providing 3.8% mAP and 1.4% EDS improvements.
+Regarding qualitative results, Figure˜5 further illustrates a case of the zero-shot capability in the real world.
+Despite being trained solely on the nuScenes dataset, TTC-MonoDETR can detect a Deer using a visual prompt of another deer from a different viewpoint (simulating an internet image search for road description “deers might appear on this road”), which is challenging for traditional offline detectors.
+More qualitative examples can be found in Appendix Appendix˜C.
 These experiments demonstrate TTC as an effective and versatile system for instant error correction, excelling at handling missing distant objects, unseen object categories, and domain shifts, without requiring any additional training.
@@
 4.3 Robustness of Visual Prompts
-As a crucial component of the continuous test-time error correction system, we conduct a series of ablation studies on visual prompts in this section.
-We base TTC system on MonoDETR and validate its robustness concerning prompts obtained from the Internet or those from different moments.
+As a crucial component of the continuous test-time error correction system, we conduct a series of experiments on visual prompts in this section.
+Specifically, we analyze the robustness of visual prompts against: (i) variations in prompt sources and styles, (ii) changes in viewpoints across frames, (iii) spatial perturbations in provided prompts, (iv) reduced feedback frequency, (v) noisy (off-topic) visual prompts, and (vi) rapidly changing prompts.
 Figure 6: Visual prompt examples derived from the Internet.
@@
 Category Method mAP (%) $\uparrow$ EDS (%) $\uparrow$
-Car MonoDETR 3.2 9.1
-TTC-MonoDETR 20.9 (+17.7) 27.5 (+18.4)
-Bus MonoDETR 0.4 4.4
-TTC-MonoDETR 17.4 (+17.0) 20.6 (+16.2)
+Car MonoDETR 3.2 37.5
+TTC-MonoDETR 20.9 (+17.7) 45.3 (+7.8)
+Bus MonoDETR 0.4 28.1
+TTC-MonoDETR 17.4 (+17.0) 40.2 (+12.1)
 Figure 7: Experiments on visual prompts from arbitrary temporal frames. TTC-3D Detectors can effectively locate and detect target objects in future frames, using its image patch at Frame #0.
 Robustness over Web-derived Visual Prompts.
-Visual prompts can be arbitrary imagery views of target objects and can be from any image source.
+Visual prompts can be arbitrary imagery views of target objects and can be from any image source and styles.
 For example, we can use images sourced from the Internet as visual prompts.
 We fix the prompt buffer with visual prompts from the Internet during inference, and assess the effectiveness of handling visual prompts with diverse styles.
-We employ the model from Table 2(b) (TTC-MonoDETR trained with 20% labels of vehicles), and select 12 car and 6 bus images from the Internet, which resemble those in nuScenes dataset, as visual prompts for online correction.
-We show them in Figure 6.
-As listed in Table 3, TTC system demonstrates strong online correction capabilities though with limited annotations.
-Even with prompts sourced from the Internet with various styles and poses, TTC system still improves the offline-trained baseline by 18.4% EDS on “Car” objects under this extremely challenging setting.
+We employ the model from Table˜II(b) (TTC-MonoDETR trained with 20% labels of vehicles), and select 12 car and 6 bus images from the Internet, which resemble those in the nuScenes dataset, as visual prompts for online correction.
+We show them in Figure˜6.
+As listed in Table˜3, TTC system demonstrates strong online correction capabilities.
+Even with prompts sourced from the Internet with various styles and poses, TTC system still improves the offline-trained baseline by 17.7% mAP on “Car” objects under this extremely challenging setting.
 This further underscores the robustness of TTC system and highlights its potential to address the long-tail challenges in real-world scenarios.
@@
 We also investigate the capability of TTC system in handling visual prompts of target objects across scenes and times.
-In Figure 7, we study whether our system can successfully associate objects with its visual prompts from arbitrary frames.
+In Figure˜7, we study whether our system can successfully associate objects with their visual prompts from arbitrary frames.
 Specifically, for each video clip of the nuScenes validation set, we use the image patch of target objects in the first frame as visual prompts, then detect and track target objects in subsequent frames, and compute the recall rate for evaluation with arbitrary views.
-Figure 7 presents the results of TTC-MonoDETR, showing that despite significant differences in viewpoints and object poses between frames, the recall rate does not drop dramatically as the ego vehicle moves.
-This highlights the robustness of the TTC detectors in handling visual prompts with arbitrary views across scenes and times, indicating that the TTC system can effectively process human feedback that may involve temporal delays.
-4.4 More Ablation Studies
-Due to limited space, we refer readers to Appendix D of the appendix for extensive ablation studies.
-These studies include: (1.) the ablations on core components in visual prompt alignment in Section D.1 and Section D.2, (2.) the evaluation of visual prompt alignment’s impact on ensuring instance awareness in Section Section D.3, (3.) the examination of statistics regarding the visual prompt buffer during online inference in Section D.4, (4.) the assessment of visual prompt robustness against user perturbations and reduced user feedback collection frequency in Section D.5 and Section D.6, (5.) the comparison highlighting the superiority of our OA module over the combination of 2D single-object trackers with 3D heads for generating 3D tracking in Section Section D.7, and more.
+Figure˜7 presents the results of TTC-MonoDETR, showing that despite significant differences in viewpoints and object poses between frames, the recall rate does not drop dramatically as the ego vehicle moves.
+This highlights the robustness of the TTC detectors in handling visual prompts with arbitrary views across scenes and times, indicating that the TTC system can effectively process feedback that may involve temporal delays.
+Robustness over Positional Perturbations.
+Since visual prompts during online inference are derived from user clicks or 2D detectors, they may deviate from the precise 2D center of the target object.
+To analyze the impact of such positional shifts, we introduce controlled perturbations during test-time inference.
+Specifically, we apply translation offsets to the ground-truth 2D centers, varying the maximum displacement ratio from 0% to 40%, and evaluate the system’s robustness under these conditions.
+As shown in Figure˜8, despite increasing perturbations, TTC-MonoDETR exhibits only minor performance degradation, demonstrating strong resilience to positional variations.
+This underscores the robustness of the TTC system in handling various interactions, where precise locations may not always be guaranteed.
+Figure 8: Robustness over positional perturbations.
+We evaluate TTC-MonoDETR under controlled perturbations during test-time inference.
+Specifically, we introduce translation offsets to the ground-truth 2D centers with maximum ratios of 0%, 10%, 20%, 30%, and 40%.
+Robustness over Visual Prompt Collection Frequency.
+In practical deployment, human feedback and open-vocabulary 2D detectors (e.g., Grounding DINO) are often delayed, making it infeasible to provide corrections for every missed detection in real time.
+To evaluate the robustness of TTC under varying feedback frequencies, we design experiments that simulate different feedback intervals.
+TABLE 4: mAP comparisons under different feedback collection intervals. “N=0” means feedback is collected every frame, while “N=2” means every 2 frames.
+Feedback Interval NN 0 2 4 6
+TTC-BEVFormer 47.8 47.5 47.3 46.9
+TTC-Sparse4Dv2 53.5 53.1 52.4 51.8
+TTC-MV2D 51.0 50.7 50.2 49.6
+As shown in Table˜4, we update the prompt buffer at different frame intervals to mimic varying levels of feedback availability.
+Despite reducing the update frequency (from every frame to every 6 frames), TTC maintains strong error correction performance, demonstrating its resilience to infrequent feedback.
+This suggests that TTC can effectively operate in real-world scenarios where continuous supervision is impractical.
+Robustness over Noisy Visual Prompts.
+To evaluate robustness against completely misleading prompts, we augment the fixed prompt buffer used in Table˜3 with two animal exemplars (one dog, one kangaroo; prompt images in Appendix Figure˜21) in addition to vehicle prompts.
+These classes are virtually absent in nuScenes imagery and thus serve as off-topic (noisy) prompts.
+We evaluate the Car category using TTC-MonoDETR from Table˜3 (trained with 20% vehicle labels), comparing runs with and without the added animal prompts.
+As shown in Table˜5, adding these off-topic prompts do not degrade accuracy, further demonstrating TTC’s robustness to noisy prompts.
+TABLE 5: Robustness to noisy visual prompts. Comparison of fixed-buffer TTC-MonoDETR with and without off-topic animal prompts (dog & kangaroo). Evaluation follow the protocol of Table˜3.
+Prompt Type mAP $\uparrow$ EDS $\uparrow$
+Vehicle 20.9 45.3
+Vehicle + Animal 20.9 45.3
+Robustness over Rapidly Changing Prompts.
+We test TTC under fast ego–motion and frequent occlusions, where prompts appear/disappear quickly.
+To emulate this regime, we uniformly sub–sample each video by 50%, effectively doubling apparent scene speed and shortening object dwell time, while keeping the 5-frame feedback interval and all thresholds/buffer rules identical to Table˜1.
+As summarized in Table˜6, the gains over the base detectors remain consistently positive (mAP/EDS and FN/frame), albeit smaller than in the full–frame setting.
+This indicates that TTC is robust to rapidly changing prompts.
+TABLE 6: Robustness to rapidly changing prompts. We choose BEVFormer as the base detector.
+Method mAP $\uparrow$ EDS $\uparrow$ FP/frame $\downarrow$ FN/frame $\downarrow$
+BEVFormer 40.2 49.6 34.9 1.9
+TTC-BEVFormer 45.9 (+5.7) 51.9 (+2.3) 35.0 1.2
+4.4 Additional Comparisons and System Analysis
+Beyond evaluating the effectiveness of TTC and the robustness of visual prompts, we conduct additional comparisons to clarify its advantages over alternative approaches: (i) comparison with an SOT-based online error–correction baseline and (ii) contrast with offline fine–tuning driven by test–time 2D feedback.
+We also provide system–level analyses: (iii) efficiency in terms of model size and inference latency; (iv) cross–dataset transfer ability; (v) feasibility of fully automated feedback based on 2D–3D misalignment; and (vi) FP trade-off induced by visual prompts.
+Comparison with SOT-based Online Error Correction.
+An intuitive way to integrate online error correction is to combine a 2D tracking algorithm with a 3D detection head.
+To explore this possibility, we implement a baseline using ATOM [danelljan2019atom] for 2D tracking and MV2D [wang2023object] as the 3D object detector.
+MV2D leverages ATOM’s tracking results of missing objects from previous frames to initialize queries, mimicking an error correction mechanism.
+In this setup, we perform human intervention every 5 frames to provide a list of missing objects along with their 2D bounding boxes for ATOM as the reference.
+For a fair comparison, the visual prompts of TTC-MV2D are derived from user-drawn bounding boxes every 5 frames.
+As shown in Table˜7, TTC, equipped with the OA module, significantly outperforms this SOT-based approach in error correction.
+This highlights the limitations of simply combining SOT with a 3D detector and underscores the necessity of a dedicated online correction framework.
+TABLE 7: Comparison of error correction performance with SOT + 3D detector head.
+The results demonstrate the superior performance of TTC in online error correction, highlighting the limitations of the baseline approach where simply combining SOT with a 3D detection head does not effectively address error correction.
+Method mAP (%) $\uparrow$ EDS (%) $\uparrow$
+MV2D 42.6 51.9
+SOT + MV2D 42.1 51.2
+TTC-MV2D 50.0 57.3
+Comparison with Offline Fine-tuning on Test-time Feedback.
+An alternative way to leverage auxiliary feedback is offline fine-tuning, where missing objects are collected during inference and later used to fine-tune the model with test-time 2D ground truth.
+While this enables adaptation, it introduces inherent delays compared to real-time correction.
+We evaluate this strategy by fine-tuning MV2D [wang2023object], which relies on 2D detection for 3D object inference and can incorporate 2D feedback.
+In this experiment, we collect human 2D feedback at different intervals and compare MV2D’s fine-tuned performance with TTC-MV2D, where prompts are updated at the same intervals.
+Results in Table˜8 show that TTC-MV2D consistently outperforms the fine-tuned MV2D with 2D weak feedback.
+This demonstrates its effectiveness of TTC in real-time streaming applications with only weak feedback feasible.
+TABLE 8: mAP comparison between TTC-MV2D and MV2D fine-tuned with 2D feedback annotations.
+“N=0” means collecting feedback at every frame, while “N=2” means every 2 frames.
+TTC-MV2D consistently outperforms the offline fine-tuned MV2D.
+Feedback Interval NN 0 2 4 6
+MV2D + Offline fine-tune 43.4 42.7 41.0 39.9
+TTC-MV2D 51.0 50.7 50.2 49.6
+Efficiency Analysis: Model Size and Latency.
+For online deployment, a system must balance accuracy and efficiency.
+We analyze TTC ’s model size and inference latency, particularly in comparison with recent LLM-based promptable 3D detection methods [chen2024ll3da, huang2023embodied].
+TABLE 9: Parameters and latency comparisons between LLM based 3D detectors, traditional 3D detectors, and related TTC 3D detectors.
+Method LL3DA MonoDETR TTC-MonoDETR
+#Params $\downarrow$ 118M 68M 80M
+FPS (Hz) $\uparrow$ 0.42 11.1 9.1
+Latency in promptable systems arises from two sources: (i) response time in providing prompts and (ii) the system’s inference speed.
+While the first factor is unavoidable, the second is crucial for real-time applications.
+As shown in Table˜9, LLM-based methods like LL3DA [chen2024ll3da] exhibit significantly higher latency, making them impractical for autonomous driving.
+In contrast, TTC introduces minimal overhead compared to its base detector while maintaining real-time performance, confirming its suitability for online deployment in safety-critical scenarios.
+Cross-dataset Transfer Analysis.
+We complement the results with a cross-dataset study from nuScenes to Waymo.
+Using MonoDETR as the base detector, Table˜10 shows that direct transfer causes a severe collapse (42.6 $\to$ 0.7 mAP).
+Accordingly, directly applying TTC brings only a marginal gain of +0.3 mAP.
+As discussed in Cross-Dataset Sensor Alignment [zheng2023cross], for vision-only 3D detectors the dominant factor is the mismatch in camera intrinsics/extrinsics: when moving from nuScenes to Waymo, objects at comparable 3D distance/shape occupy markedly different 2D pixel scales due to focal-length disparities and rig configurations (Appendix Figure˜11), which induces depth bias.
+To decouple this confound, the intrinsic synchronization (K-sync) strategy [zheng2023cross] is adopted: images are resized and cropped to normalize focal lengths prior to inference.
+Under K-sync, TTC yields a substantial improvement over the K-sync–only baseline on Waymo ( +3.3 mAP ), indicating that once intrinsics are aligned, the proposed prediction-level rectification generalizes across datasets without updating model weights (Table˜10).
+TABLE 10: mAP results of cross-dataset validation. The model is trained on nuScenes and evaluated on the nuScenes and Waymo. MonoDETR is used as the base detector.
+Method nuScenes Waymo
+MonoDETR 42.6 0.7
+TTC-MonoDETR 45.0 1.0
+K-sync + MonoDETR 42.8 30.8
+K-sync + TTC-MonoDETR 45.8 (+3.0) 34.1 (+3.3)
+Fully Automated Feedback Analysis.
+To address the practical difficulty of obtaining clicks/boxes while driving, we evaluate a most realistic deployment mode in which all prompts are generated automatically (no human inputs).
+We harvest feedback from the discrepancy between a 2D detector and the image reprojection of the current 3D predictions.
+Concretely, we consider two 2D detectors: (i) off-the-shelf Grounding DINO [liu2024grounding], and (ii) a YOLOv8 [ultralytics_yolov8] model retrained on the nuScenes 2D taxonomy.
+For each frame, every 2D box is compared against the projected 3D boxes; unmatched cases (IoU <0.5<0.5; optionally gated by reprojection residual) are converted into instance-specific visual prompts by cropping the corresponding image patch and enqueuing it into the buffer.
+Under the same in-domain protocol as Table˜1, automated prompts yield consistent improvements in mAP/EDS (Table˜11).
+TABLE 11: Automated feedback results.
+2D–3D misalignment is computed between 2D boxes (Grounding DINO / YOLOv8-nuScenes) and the image reprojection of 3D predictions; unmatched 2D boxes are converted into visual prompts.
+Method mAP $\uparrow$ EDS $\uparrow$
+BEVFormer 41.1 50.2
+TTC-BEVFormer (GDINO) 41.8 (+0.7) 50.4 (+0.2)
+TTC-BEVFormer (YOLOv8) 42.4 (+1.3) 50.8 (+0.6)
+FP Impact Analysis.
+Because visual prompts introduce additional hypotheses at inference, a natural concern is whether they inflate false alarms.
+We therefore quantify the effect of TTC on false positives in Table˜1.
+Concretely, we report FP/frame and FN/frame over the full validation stream, where $\text{FP/frame}=\tfrac{\text{FP}_{\text{total}}}{\text{frames}}$ and $\text{FN/frame}=\tfrac{\text{FN}_{\text{total}}}{\text{frames}}$.
+A detection is matched using the same center–distance protocol as mAP calculation (averaged over $\{0.5,1.0,2.0,4.0\}$ m).
+Across diverse backbones, TTC markedly reduces FN/frame, while FP/frame changes only marginally ($\leq\!+0.2$), indicating that the recall gains do not come at the cost of inflating false alarms.
+We also evaluate a deployment-oriented setting with a confidence threshold of $0.3$; the FP/frame is not statistically significant before vs.
+after applying TTC (see Appendix Table˜15).
+Why FP remains controlled.
+As stated in our method, each prompt spawns up to $N$ hypotheses (N=4N{=}4).
+For clarity, we define a valid prompt is one whose associated object is present in the current scene, whereas an invalid prompt is stale or contradictory (the object has left the scene or has no corresponding instance).
+On TTC–BEVFormer, we observe on average only $3.9$ visual prompts per frame; each valid prompt yields 1.6 above-threshold candidates on average, and each invalid prompt yields 0.3, indicating that visual prompts do not flood the detector and that invalid prompts are usually low-confidence.
+For above-threshold candidates, the outcomes are: (i) correct match to the intended object (desired); (ii) match to a visually similar object that has already been detected—suppressed by NMS, so no extra box appears; (iii) match to a similar object that is not detected—this becomes a TP for that object rather than an FP; (iv) fall on an empty region but overlap an FP from the dense query results—removed by NMS, thus not affecting the final FP count; (v) fall on an empty region with no overlap—this is a true FP, but it is rare (0.2 per frame in our statistics).
+In practice, (i) and (ii) dominate, while (v) is negligible; consequently, FP/frame stays near baseline as FN/frame drops (Appendix Figure˜12).
+4.5 Ablation Studies
+To better understand the key design choices in TTC, we conduct a series of ablation studies.
+Specifically, we investigate: (i) the contribution of core alignment components such as similarity and localization losses, (ii) the effect of addressing instance ambiguity through a one-to-$N$ mapping strategy, (iii) the capability of visual prompt alignment in achieving instance-level object differentiation, and (iv) the dynamic behavior of the visual prompt buffer during online inference.
+These studies provide deeper insights into the effectiveness and robustness of TTC in leveraging visual prompts for online 3D detection.
+Effect of Components in Visual Prompt Alignment.
+TABLE 12: Effect of Visual Prompt Alignment. We select TTC-MonoDETR for this experiment.
+Sim. Loss Loc. Loss mAP (%) EDS (%)
+- - 32.6 44.2
+$\surd$ - 39.4 48.5
+$\surd$ $\surd$ 43.3 52.3
+Visual prompt alignment aims to localize objects via visual prompts in input images.
+We now evaluate its components with TTC-MonoDETR.
+Table˜12 presents experiments to verify the core components of this alignment, including similarity loss (Focal and Dice loss) and localization loss (supervising visual prompt localization, $\mathcal{X}_{v}$).
+While the alignment can be implicitly learned by attention mechanisms in the transformer decoder, incorporating explicit similarity supervision brings improvements of 6.8% mAP and 4.3% EDS.
+Further utilizing the position loss and one-to-$N$ mapping (to address instance ambiguity) boosts the performance to 43.3% mAP and 52.3% EDS.
+These results prove this alignment operation is a critical component enabling the TTC detectors to effectively detect target objects via visual prompts.
+Figure 9: Visual prompt showcase for Table˜13.
+TABLE 13: Effect of visual prompt alignment on instance awareness.
+Group (a.) resembles “car” objects in nuScenes validation set with diverse types and colors; while group (b.) maintains black sedans only.
+The performance gap demonstrates that alignment learns the instance awareness.
+Group mAP (%) $\uparrow$ EDS (%) $\uparrow$
+a. 20.9 45.3
+b. 13.8 42.2
+Instance Ambiguity in Visual Prompt Alignment.
+TABLE 14: Effect of the number of predicted positions $N$ of visual prompt alignment. We select TTC-MonoDETR for this experiment.
+No. of Position Predictions mAP (%) $\uparrow$ Recall (%) $\uparrow$
+1 39.9 62.1
+4 43.3 69.1
+8 43.0 69.4
+To solve the instance ambiguity issue, we propose to regress $N$ positions of each visual prompt when performing the visual prompt alignment.
+This retrieves all objects with similar visual contents.
+In this study, we validate the effectiveness of this design by conducting ablation studies on the number of $N$.
+As listed in Table˜14, when the number of position predictions $N$ equals 1, which means a one-to-one mapping for each visual prompt, the mAP and recall rate are 39.9% and 62.1%, respectively.
+Then, if we increase the $N$ to 4, effectively a one-to-four mapping, we obtain an mAP of 43.3% and a recall rate of 69.1%.
+This represents a 7% improvement in the recall rate, demonstrating that instance ambiguity is an important challenge in visual prompt-based detection, and the proposed one-to-$N$ mapping solution effectively addresses this issue.
+Effect of Visual Prompt Alignment on Instance Awareness.
+Despite demonstrating that visual prompt alignment can enhance system performance, we remain uncertain whether the alignment can distinguish different objects based on visual prompts for instance-level matching.
+Figure 10: Size of visual prompt buffer during the video stream.
+Visual prompt buffer stores the missed objects during online inference to rectify test-time errors of deployed 3D detectors.
+It can adaptively manage the stored prompts and thus maintain the balance between latency and accuracy.
+To investigate, we conduct an additional experiment, based on a similar setting with Table˜3 (TTC-MonoDETR trained on 20% vehicle annotations and freeze the prompt buffer with predefined web prompts when inference.), but fix the visual prompt buffer with images of black sedans solely (Figure˜9).
+As shown in Table˜13, as “Car” objects in nuScenes contain various types and colors, solely using black sedans as visual prompts leads to a 7.1% mAP drop.
+This underscores that the alignment operation effectively differentiates objects based on visual prompts, achieving instance-level matching and detection.
+Statistics of Visual Prompt Buffer during Online Inference.
+We design the visual prompt buffer to store missed objects during inference with the video stream and introduce a “dequeue” mechanism to prevent the buffer from growing indefinitely.
+In this ablation study, we analyze the dynamic buffer size, as well as the number of enqueued and dequeued in each frame, to illustrate the behavior of the visual prompt buffer during the online operation of TTC.
+As shown in Figure˜10, the visual prompt buffer exhibits three distinct behaviors during online inference in each nuScenes video clip: increasing, steady, and decreasing.
+In initial frames, many traffic cones are queued into the buffer due to the poor performance of deployed offline 3D detectors on cone objects.
+The buffer size thus grows quickly in initial frames to store visual prompts of missed objects for online rectification (Frames #0 to #4).
+The buffer size then stabilizes as the online detector consistently detects and tracks all objects of interest (Frames #4 to #20).
+Further, as the ego vehicle drives out of the scene, many previously enqueued objects no longer exist and are thus removed from the buffer automatically (Frames #20 to #40).
+The buffer finally becomes empty as the scene changes.
+This demonstrates the effectiveness of the visual prompt buffer, which consistently stores missed objects during online inference and corrects online errors.
+This dynamic behavior, exhibiting increasing, steady, and decreasing phases, highlights itself to manage stored visual prompts for robust 3D object detection performance throughout the online inference with the balance between latency and accuracy.
+We also report several compact, deployment-oriented statistics over the full validation stream on TTC-BEVFormer: (i) the largest buffer size observed over the entire stream is only 65, with an average of 3.9; (ii) after an object leaves the scene, only 10.3% of its prompts remain in the buffer; (iii) only 17.2% of those left objects in (ii) produce erroneous final box after standard post-processing (confidence filtering + NMS).
+Together, these results indicate that bounded buffer growth, timely removal of outdated prompts, and negligible impact on final predictions—consistent with the dynamics in Figure˜10.
 5 Conclusion
@@
 For performance, the generalization ability is constrained by the model size and data volume, and future work will aim to enhance this by scaling up both the data and parameters.
-For model design, while allowing user interaction provides reliability, it introduces potential risks.
-Requiring frequent human feedback may be a challenge in current L3 autonomous driving paradigms.
-Though reducing the frequency of feedback can alleviate this pressure without compromising performance (See appendix Section D.6), there is room for improvement.
+For model design, while allowing auxiliary information to provide reliability, it introduces potential risks.
+Requiring frequent feedback, especially human feedback, may be challenging in current L3 autonomous driving paradigms.
+Though reducing the frequency of feedback can alleviate this pressure without compromising performance, there is room for improvement.
 Future work could integrate robust tracking methods or leverage advanced hardware, such as eye-tracking, to simplify the feedback collection.
@@
 Acknowledgement
-OpenDriveLab is the autonomous driving team affiliated with Shanghai AI Lab.
 This work was supported by National Key R&D Program of China (2022ZD0160104), NSFC (62206172), and Shanghai Committee of Science and Technology (23YF1462000).
+This work is in part supported by the JC STEM Lab of Autonomous Intelligent Systems funded by The Hong Kong Jockey Club Charities Trust.
 We thank team members from OpenDriveLab for valuable feedback along the project.
 Special thanks to Chengen Xie for providing data support, and Jiahui Fu for sharing insights on MV2D.
+We also sincerely thank Qingsong Yao for providing writing guidance and suggestions on experimental design.
 Appendix A Discussions
@@
 We emphasize, in this paper, we do not propose the online system to replace the traditional offline data loop.
-As illustrated in Fig.
-2 of the main paper, these two systems address different aspects of autonomous driving.
+As illustrated in Figure˜2, these two systems address different aspects of autonomous driving.
 The offline system remains crucial for enhancing the capabilities of the base perception model through development; while our online TTC system further enables the deployed frozen model in vehicles to promptly rectify dangerous driving behaviors caused by unrecognized objects on the road.
-With improved offline-trained detectors, the TTC can effectively correct more online errors during test-time inference without re-training, as detailed in Table 1 of the main paper.
+With improved offline-trained detectors, the TTC can effectively correct more online errors during test-time inference without re-training, as detailed in Table˜1.
 Q2: What are the main technical novelty and advantages of the proposed TTC system over previous instruction-based 3D detectors?
@@
 In contrast, verbal descriptions can be ambiguous to convey instance-level features, leading to an inaccurate understanding of the missing objects.
-Second, text promptable models are often combined with LLMs with high latency and are thus unavailable for autonomous driving deployment, as discussed in Section D.8.
+Second, text promptable models are often combined with LLMs with high latency and are thus unavailable for autonomous driving deployment, as discussed in Section˜4.4.
 Box and point prompts are less convenient than visual prompts when dealing with stream data.
-If missing occurs, these single-frame prompts require users to provide feedback at every frame, which is unfeasible in real-world applications.
+If missing occurs, these single-frame prompts require feedback at every frame, which is unfeasible in real-world applications.
 Compared to box and point prompts, visual prompts are robust across different scenes and timestamps, one single-frame visual prompt is enough for detecting and tracking in later frames.
-Furthermore, visual prompts enable 3D detection with pre-defined visual descriptions of target objects, regardless of the sources, styles, poses, etc, as discussed in Section 4.3.
-The introduction of novel visual prompts enables real-time, accurate, and continuous error correction of streaming inputs with ”one-click” feedback.
+Furthermore, visual prompts enable 3D detection with pre-defined visual descriptions of target objects, regardless of the sources, styles, poses, etc, as discussed in Section˜4.3.
+The introduction of novel visual prompts enables real-time, accurate, and continuous error correction of streaming inputs.
 Q3: What is the relationship between the TTC system and existing 3D perception tasks?
@@
 In summary, the TTC system represents a more flexible and comprehensive 3D object detection framework, combining the strengths of zero-shot detection, handling out-of-distribution objects, and utilizing diverse visual prompts beyond the current scene context.
-Q4: Why choose 3D detection as the experimental scenarios of TTC? Could the proposed framework be extended to 2D detection or other vision tasks?
-TTC represents a general idea to equip deployed systems with the capability of online error rectification, making them more versatile, adaptive, and reliable.
-This idea can be readily extended to other vision tasks, such as 2D detection, for rapid adaptation of pre-trained models to novel scenarios.
-The choice of 3D detection for autonomous driving as the experimental scenario is motivated by the paramount importance of safety for deployed self-driving systems.
-Without an online correction method, mistakes made by the offline model pose significant safety risks for on-road autonomy.
-Therefore, we select the autonomous driving domain as the testbed for the TTC framework, given the critical need for a robust, adaptive online error correction system to ensure the reliability of these safety-critical applications.
-We mark the extension of TTC to other vision tasks as future works.
-Q5: What are potential applications and future directions of TTC?
+Q4: What are potential applications and future directions of TTC?
 We believe that, visual prompts, as the core design element of the TTC system, represent a more natural and intuitive query modality for the image domain.
@@
 As showcased in this work, the visual prompt-based framework opens up new possibilities for online perception systems, not only in autonomous driving but also in a broader range of domains.
-Appendix B More Related Work
-B.1 Interactive Vision Models.
-Interactive vision models are designed for tasks based on user inputs.
-As one of the fundamental tasks in computer vision, they are extensively researched with numerous breakthroughs [26, 6, 33, 60, 14].
-Particularly, the advent of the Segment Anything Model (SAM) [24] has sparked a surge of progress, with applications spanning reconstruction [49], detection [47, 63], segmentation [71], image editing [13], and more.
-Compared to existing models, which typically rely on prompts such as clicks, boxes, or scribbles, in this work, we study visual prompts, a new prompt referring to the actual images of objects with arbitrary poses and styles.
-Visual prompts enable continuous detection and tracking of target objects, facilitating the immediate correction of failed detections at test time.
-This represents a departure from traditional prompt types, offering more natural and dynamic interactions with the visual content.
-B.2 Test-time Adaptation
-Test-time adaptation aims to improve model performance on test data by adapting the model to test samples, even in the presence of data shifts [69, 38, 39, 48, 53, 34].
-However, this area remains relatively unexplored in the context of 3D detection.
-Recent work, such as MonoTTA [29], has explored domain adaptation for 3D detection by fine-tuning models on batch streaming data under different weather conditions.
-In contrast, online 3D detection does not modify the original model parameters.
-Instead, it leverages human feedback to achieve instance-level error correction, thereby enabling domain adaptation without compromising the model’s performance on the original domain.
-Appendix C Implementation Details
-C.1 Details of TTC System
-Figure 8: Details of the TTC system.
-This section elaborates on the detailed workflow of the TTC system.
-As shown in Figure 8, during online inference, once the TTC-3D Detector fails to recognize an object, users can click on the missing object within the image.
-Based on the user-provided click, the TTC-3D Detector identifies the corresponding 2D boxes, crops the relevant areas to obtain visual prompt patches, and subsequently updates the visual prompt buffer.
-In later frames of inference, the TTC-3D Detector applies these stored visual prompts to continuously detect and track previously missing 3D objects, achieving instant error correction during test-time and constantly enhancing the offline-trained 3D detectors after being deployed.
-C.2 Visual Prompt Alignment
-Figure 9: Concrete mechanism of visual prompt alignment.
-This figure illustrates monocular input.
-When multi-view images are employed, this alignment operation flattens the different views and still generates $N$ peak candidate positions.
-We now delve into the implementation details of visual prompt alignment.
-As described in Figure 9, given the visual prompt $\mathcal{P}_{v}$, we first extract the visual prompt features $\mathcal{Z}_{v}$ using a lightweight encoder, e.g., ResNet18 [17] in our implementation.
-Then, we use two separate 2-layer perceptrons, each with 128 and 64 output channels, to align the channel dimensions of the prompt features and the input image features.
-This results in the aligned feature maps $\hat{\mathcal{Z}_{v}}\in\mathbb{R}^{M\times 64}$ and $\hat{\mathcal{Z}}\in\mathbb{R}^{KHW\times 64}$, where $M$ is the number of visual prompts, $K$ is the number of image views, $H$ and $W$ are the spatial dimensions of image features.
-Finally, we compute the cosine similarity ($\odot$ in Figure 9) between $\hat{\mathcal{Z}}_{v}$ and $\hat{\mathcal{Z}}$ to obtain the similarity map $\mathcal{S}\in\mathbb{R}^{M\times KHW}$, which encodes the alignments between the visual prompts and image features.
-To solve the instance ambiguity issue, we propose to predict multiple spatial coordinates of different peak responses in the similarity map (highlighted by “orange box” in Figure 9).
-Specifically, we first multiply the image features $\mathcal{Z}$ with similarity map $\mathcal{S}$ ($\bigotimes$ in Figure 9), and then use a 2-layer perceptron with $N\times 2$ output channels to regress spatial coordinates of the $N$ peak responses.
-During training, we only backpropagate the localization loss for the positions that have the minimum loss with respect to the ground truth.
-In our implementation, $N$ equals 4.
-This multi-instance retrieval approach allows the TTC to handle cases where the visual prompt matches multiple candidate objects in the input image, improving the robustness of the online error correction.
-C.3 Details of Entity Detection Score (EDS)
-Table 4: Feasibility Analysis of the EDS Metric.
-Method mAP (%) w.o. Rec (%) + Rec (%)
-BEVFormer 36.5 49.1 35.8
-TTC-BEVFormer 47.8 54.6 47.2
-BEVFormer-V2 39.6 49.2 38.9
-TTC-BEVFormer-V2 51.6 58.5 51.0
-EDS is a class-agnostic version of nuScenes Detection Score (NDS), but further prioritizes the localization quality of target objects.
-Inspired by [64], we improve the original NDS computation by multiplying the recall rate and the mean True Positive metrics, $\mathbb{TP}$, as well, as illustrated below:
-$\text{EDS}\!=\!\frac{1}{6}[3\text{mAP}\!+\!\text{Recall}\!\times\!\sum_{\text{mTP}\in\mathbb{TP}}(1-\text{min}(1,\text{mTP}))],$ (2)
-The intuition behind this is simple.
-The larger the recall rate is, the more predictions are involved in the statistics of mTP.
-Compared to simply setting a recall threshold [1], the multiplication adjusts the weight of mTP to EDS according to its comprehensiveness and thus brings a more informative quantitative result.
-In Table 4, we analyze the effectiveness of this metric.
-The comparisons between multiplying recall rate or not on various TTC-3D Detectors show that EDS, incorporating recall into its calculation, does not alter the original overall trend.
-Furthermore, it effectively highlights the superiority of detecting missed objects, demonstrating its validity.
-Appendix D Ablation Studies
-In this section, we conduct a series of ablation studies.
-We base our TTC system on MonoDETR for its simple and efficient monocular setting except for experiments in Section D.6, Section D.7 and Section D.9.
-D.1 Effect of Components in Visual Prompt Alignment.
-Table 5: Effect of visual prompt alignment.
-Sim. Loss Loc. Loss mAP (%) EDS (%)
-- - 32.6 31.9
-$\surd$ - 39.4 38.5
-$\surd$ $\surd$ 43.3 42.8
-Visual prompt alignment aims to localize objects via visual prompts in input images.
-We now evaluate its components.
-Table 5 presents experiments to verify the core components of this alignment, including similarity loss (Focal and Dice loss) and localization loss (supervising visual prompt localization, $\mathcal{X}_{v}$).
-While the alignment can be implicitly learned by attention mechanisms in the transformer decoder, incorporating explicit similarity supervision brings improvements of 6.8% mAP and 6.6% EDS.
-Further utilizing the position loss and one-to-$N$ mapping (to address instance ambiguity) boosts the performance to 43.3% mAP and 42.8% EDS.
-These results prove this alignment operation is a critical component enabling the TTC detectors to effectively detect target objects via visual prompts.
-D.2 Instance Ambiguity in Visual Prompt Alignment
-Table 6: Effect of the number of predicted positions $N$ of visual prompt alignment.
-No. of Position Predictions mAP (%) $\uparrow$ Recall (%) $\uparrow$
-1 39.9 62.1
-4 43.3 69.1
-8 43.0 69.4
-To solve the instance ambiguity issue, we propose to regress $N$ positions of each visual prompt when performing the visual prompt alignment.
-This retrieves all objects with similar visual contents.
-In this study, we validate the effectiveness of this design by conducting ablation studies on the number of $N$.
-As listed in Table 6, when the number of position prediction $N$ equals 1, which means a one-to-one mapping for each visual prompt, the mAP and recall rate are 39.9% and 62.1%, respectively.
-Then, if we increase the $N$ to 4, effectively a one-to-four mapping, we obtain an mAP of 43.3% and a recall rate of 69.1%.
-This represents a 7% improvement in the recall rate, demonstrating that instance ambiguity is an important challenge in visual prompt-based detection, and the proposed one-to-$N$ mapping solution effectively addresses this issue.
-D.3 Effect of Visual Prompt Alignment on Instance Awareness
-Figure 10: Size of visual prompt buffer during the video stream.
-Visual prompt buffer stores the missed objects during online inference to rectify test-time errors of deployed 3D detectors.
-It can adaptively manage the stored prompts and thus maintain the balance between latency and accuracy.
-Figure 11: Visual prompt showcase for Table 7.
-Table 7: Effect of visual prompt alignment on instance awareness.
-Group (a.) resembles “car” objects in nuScenes validation set with diverse types and colors; while group (b.) maintains black sedans only.
-The performance gap demonstrates that alignment learns the instance awareness.
-Group mAP (%) $\uparrow$ EDS (%) $\uparrow$
-a. 20.9 27.5
-b. 13.8 18.9
-Despite demonstrating that visual prompt alignment can enhance system performance, we remain uncertain whether the alignment can distinguish different objects based on visual prompts for instance-level matching.
-To investigate, we conduct an additional experiment, based on a similar setting with Table 3 of the main paper [^3] For reference, we employ the TTC-MonoDETR trained on 20% vehicle annotations and freeze the prompt buffer with predefined web prompts when inference., but fix the visual prompt buffer with images of black sedans solely (Figure 11).
-As shown in Table 7, as “Car” objects in nuScenes contain various types and colors, solely using black sedans as visual prompts leads to an 8.6% EDS drop.
-This underscores that the alignment operation effectively differentiates objects based on visual prompts, achieving instance-level matching and detection.
-D.4 Statistics of Visual Prompt Buffer during Online Inference.
-We design the visual prompt buffer to store missed objects during inference with video stream and introduce a “dequeue” mechanism to prevent the buffer from growing indefinitely.
-In this ablation study, we analyze the dynamic buffer size, as well as the number of enqueued and dequeued in each frame, to illustrate the behavior of visual prompt buffer during the online operation of TTC.
-As shown in Figure 10, the visual prompt buffer exhibits three distinct behaviors during online inference in each nuScenes video clip: increasing, steady, and decreasing.
-In initial frames, many traffic cones are queued into the buffer due to the poor performance of deployed offline 3D detectors on cone objects.
-The buffer size thus grows quickly in initial frames to store visual prompts of missed objects for online rectification (Frames #0 to #4).
-The buffer size then stabilizes as the online detector consistently detects and tracks all objects of interest (Frames #4 to #20).
-Further, as the ego vehicle drives out of the scene, many previously enqueued objects no longer exist and are thus removed from the buffer automatically (Frames #20 to #40).
-The buffer finally becomes empty as the scene changes.
-This demonstrates the effectiveness of visual prompt buffer, which consistently stores missed objects during online inference and corrects online errors.
-This dynamic behavior, exhibiting increasing, steady, and decreasing phases, highlights itself to manage stored visual prompts for robust 3D object detection performance throughout the online inference with the balance between latency and accuracy.
-D.5 Visual Prompts with User Pertubations
-Figure 12: Robustness against feedback noise.
-In Section 4.3, we have demonstrated the robustness of our system on visual prompts from diverse sources, styles, poses, scenes, and timestamps.
-Considering visual prompts during online inference are derived from user clicking, this will have positional deviations to the perfect 2D center of the target object.
-Thus, we design experiments with positional perturbations during test-time inference to analyze the impact.
-Expressly, we set the maximum translation ratios at 0%, 10%, 20%, 30%, and 40% to the ground-truth 2D centers, and input them to TTC system to evaluate the robustness against such disturbances.
-Figure 12 shows that the TTC-MonoDETR exhibits minor performance drops under increasing positional perturbations, demonstrating strong robustness to these disturbances and highlighting significant potential for real-world applications.
-D.6 Robustness of TTC to Reduced Feedback Frequency
-Table 8: mAP comparisons of different prompt collection frequencies.
-“N=0” means having 2D feedbacks of missed detections at every frame; “N=2” means less 2D feedbacks collected every 2 frames.
-TTC system demonstrates robust error correction performance even with reduced feedback frequency.
-Exp. on Human Feedback collected with different frame interval $N$. N
-0 2 4 6
-TTC-BEVFormer 47.8 47.5 47.3 46.9
-TTC-Sparse4Dv2 53.5 53.1 52.4 51.8
-TTC-MV2D 51.0 50.7 50.2 49.6
-Considering the practical challenges of deploying this system, where human feedback inevitably involves some delay, it is difficult to provide feedback on missed detections for every frame without additional aids.
-To address this, we design experiments to simulate such scenarios and investigate the robustness of visual prompts under reduced human feedback frequency.
-As shown in Table 8, we update the prompt buffer at varying frame intervals to mimic feedback provided at different time intervals.
-TTC demonstrates robust error correction performance even with reduced feedback frequency (from every frame to every 6 frames), showcasing its feasibility for real-world applications where real-time feedback may not always be practical.
-D.7 Comparisons to Single Object Tracking (SOT) with 3D Detection Head
-Table 9: Comparison of error correction performance with SOT + 3D detector head.
-The results demonstrate the superior performance of TTC in online error correction, highlighting the limitations of the baseline approach where simply combining SOT with a 3D detection head does not effectively address error correction.
-Method mAP (%) $\uparrow$ EDS (%) $\uparrow$
-SOT + MV2D 43.0 43.2
-TTC-MV2D 51.0 51.0
-In this section, we compare our TTC system with a simple implementation of online error correction that integrates SOT and 3D detection.
-Due to the lack of related work on single object 3D tracking, we use the ATOM algorithm for 2D tracking [12] and the MV2D detector as the 3D object detection head [57].
-MV2D can leverage the 2D tracking results obtained by ATOM to initialize the query, thereby achieving the error correction ability similar to TTC system.
-For implementation, we record the missed objects of MV2D and store their reference images along with 2D bounding boxes in the prompt buffer, which are inputs ATOM requires.
-We do not perform dequeue operations to keep tracking all missing objects by ATOM.
-As shown in Table 9, the TTC system, empowered by the OA module, significantly outperforms the combination of SOT and the 3D detector head in error correction performance.
-This result highlights that simply combining SOT with a 3D detection head does not effectively achieve online error correction, further demonstrating the effectiveness of our TTC system.
-D.8 Analysis of Model Size and Latency
-Table 10: Parameters and latency comparisons between LLM based 3D detectors, traditional 3D detectors, and related TTC 3D detectors.
-Method LL3DA MonoDETR TTC-MonoDETR
-#Params 118M 68M 80M
-FPS (Hz) 0.42 11.1 9.1
-We further analyze the model size and latency of TTC system, primarily comparing them with recent LLM-based promptable 3D detection methods [4, 22].
-We argue that, for online promptable systems that are developed for real-world applications, latency can come from into two parts.
-The first one is the unavoidable delay caused by humans from observing the error to reacting to provide prompts.
-This is inherent to any online prompt-based approach.
-The other one is the delay associated with the inference speed of the system itself.
-As the first one cannot be controlled by the system design, we focus on the latter here.
-As shown in Table 10, LLM-based promtable methods, like LL3DA [4] exhibit high latency that is inadequate for autonomous driving deployments.
-In contrast, our TTC system introduces only a little extra latency compared to its base detector, which meets the real-time inference requirements and thus can be applicable for online autonomous driving systems.
-D.9 Comparison with Offline Fine-tuning using User Feedback
-Table 11: Comparisons between the TTC-MV2D and MV2D fine-tuned with feedback 2D annotations.
-“N=0” means having 2D feedbacks at every frame; “N=2” means less 2D feedbacks collected every 2 frames.
-TTC system, though without any extra training, still outperforms the offline 2D fine-tuned MV2D, especially when annotations are limited (N=6).
-Exp. on Human Feedback collected with different frame interval $N$. N
-0 2 4 6
-MV2D + Offline fine-tune 43.4 42.7 41.0 39.9
-TTC-MV2D 51.0 50.7 50.2 49.6
-In this section, we compare our TTC method with the approach that collects missing objects during inference and subsequently fine-tunes with the collected test-time 2D ground truth.
-This is another approach for utilizing test-time human feedback, though with delays in further model fine-tuning.
-We select MV2D [57] as the baseline since it relies on 2D detection results for 3D object detection, thus allowing it to utilize 2D human feedback annotations to fine-tune.
-Additionally, we collect 2D feedback from various frame intervals, as users cannot provide feedback at every frame.
-For MV2D, we use all the 2D box annotations collected at specified intervals for fine-tuning.
-For the TTC-MV2D, we update the prompt buffer at these specified intervals.
-As shown in Table 11, TTC-MV2D outperforms the MV2D model fine-tuned with 2D feedback.
-Notably, the performance of MV2D fine-tuned with human feedback from larger frame intervals, which means less frequent human feedback, declines significantly compared to models fine-tuned with feedback from every frame.
-In contrast, the TTC system allows the MV2D to perform effectively even with less frequent prompts of missed detections.
-This finding highlights the practicality of the TTC design in utilizing a single corresponding visual prompt for streaming data, as it is impractical for users to provide 2D prompts at each frame.
-Appendix E Qualitative Results
+Appendix B Additional Explanations
+This section provides visual explanations and additional tables supporting analyses in the main paper.
+Figure 11: Intrinsic mismatch across datasets.
+Equal-size image patches around cars at comparable 3D distance/shape from nuScenes (left) and Waymo (right).
+Despite similar geometry, the apparent 2D scale differs markedly due to different focal lengths and camera rigs, inducing depth bias and explaining why naive cross-dataset transfer of vision-only 3D detectors degrades severely.
+Adapted from [zheng2023cross].
+Figure 12: Outcomes of visual prompts and their FP impact.
+Final predictions consist of dense-query results and visual-prompt results.
+Visual prompts produce few above-threshold candidates, and those candidates fall into five cases (i–v).
+After confidence filtering and NMS, only a small portion contributes to false positives (v), explaining why FP/frame remains near baseline.
+B.1 Cross-dataset Transfer.
+Figure˜11 visualizes the intrinsic mismatch across datasets, adapted from [zheng2023cross].
+Equal-size image patches from nuScenes (left) and Waymo (right) show that objects at comparable 3D distance and shape appear at different 2D scales due to camera focal-length and rig disparities.
+This visualization supports the cross-dataset transfer analysis in Section˜4.4, explaining why naive cross-dataset transfer of vision-only 3D detectors leads to depth bias and severe degradation.
+B.2 False-Positive Impact.
+Figure˜12 illustrates five cases of visual prompt results.
+After confidence filtering and NMS, valid prompts mainly produce correct detections or re-detections, while invalid prompts rarely survive.
+The remaining above-threshold cases (i–v) explain why TTC reduces missed detections without materially increasing false positives, as discussed in FP Analysis in Section˜4.4.
+To corroborate the deployment-relevant regime, Table 15 reports FP/frame and FN/frame at a confidence threshold of $\tau{=}0.3$; changes in FP/frame are negligible and not statistically significant across backbones.
+TABLE 15: False-alarm metrics at a deployment-oriented confidence threshold ($\tau{=}0.3$).
+Method FP/frame@0.3 $\downarrow$ FN/frame@0.3 $\downarrow$
+MonoDETR 1.5 2.5
+TTC-MonoDETR 1.5 2.1
+MV2D 1.3 2.9
+TTC-MV2D 1.3 2.2
+Sparse4Dv2 1.1 3.0
+TTC-Sparse4Dv2 1.1 2.2
+BEVFormer 1.2 2.9
+TTC-BEVFormer 1.2 2.1
+BEVFormerV2-t8 1.2 2.9
+TTC-BEVFormerV2-t8 1.2 2.1
+RayDN 1.3 2.9
+TTC-RayDN 1.3 2.1
+StreamPETR 1.3 2.8
+TTC-StreamPETR 1.3 2.3
+Appendix C Qualitative Results
 We provide extensive visualizations to demonstrate the versatility of the TTC system across diverse scenarios:
-  • In Section E.1 and Section E.2, we fix the prompt buffer with visual prompts of either labeled or novel, unlabeled 3D objects from the nuScenes dataset to detect targets in the nuScenes images.
-  • In Section E.3, we test the performance with prompt buffer containing visual prompts in styles differing from the training distribution, such as Lego.
-  • In Section E.4, we visualize the similarity maps on out-of-domain images, including YouTube driving videos and Internet-sourced visual prompts, demonstrating the generalization of the visual prompt alignment and the effectiveness of our TTC in reducing driving risks in non-standard scenarios.
+  • In Section˜C.1 and Section˜C.2, we fix the prompt buffer with visual prompts of either labeled or novel, unlabeled 3D objects from the nuScenes dataset to detect targets in the nuScenes images.
+  • In Section˜C.3, we test the performance with prompt buffer containing visual prompts in styles differing from the training distribution, such as Lego.
+  • In Section˜C.4, we visualize the similarity maps on out-of-domain images, including YouTube driving videos and Internet-sourced visual prompts, demonstrating the generalization of the visual prompt alignment and the effectiveness of our TTC in reducing driving risks in non-standard scenarios.
 For all visualizations, the fixed prompt buffer is shown in the first row.
@@
 All results are conducted with TTC-MonoDETR.
-E.1 In-domain Visual Prompts on nuScenes “Seen” Objects
+C.1 In-domain Visual Prompts on nuScenes “Seen” Objects.
 This visualization focuses on the in-domain detection performance of the TTC system on the nuScenes dataset.
-We utilize visual prompts from labeled objects in the nuScenes dataset, and demonstrate the system’s ability to effectively detect and track these target objects across different frames, as shown in Figure 13 and Figure 14.
-The results illustrate that our TTC can accurately localize and consistently track the target objects of interest within the nuScenes scenarios, showcasing its effectiveness in handling in-domain visual prompts.
-E.2 In-domain Visual Prompts on nuScenes “Unseen” Objects
+We utilize visual prompts from labeled objects in the nuScenes dataset, and demonstrate the system’s ability to effectively detect and track these target objects across different frames, as shown in Figure˜13 and Figure˜14.
+C.2 In-domain Visual Prompts on nuScenes “Unseen” Objects
 This visualization focuses on the TTC’s ability to handle visual prompts of objects not labeled in the nuScenes dataset.
-As shown in Figure 15, Figure 16, and Figure 17, our method demonstrates its potential to detect and track novel, out-of-distribution objects with these unseen visual prompts.
-The results illustrate the TTC system’s capability to go beyond the training distribution and effectively localize and track objects that were not part of the original labeled dataset.
-This showcases the versatility and generalization ability of the visual prompt-based framework, enabling the detection of previously unseen objects.
-These findings highlight the potential of the TTC system to continuously expand its object detection capabilities by incorporating user-provided visual prompts, even for objects that were not included in the initial training data.
-E.3 Out-domain Visual Prompts on nuScenes “Seen” Objects
+As shown in Figure˜15, Figure˜16, and Figure˜17, our method demonstrates its potential to detect and track novel, out-of-distribution objects with these unseen visual prompts.
+C.3 Out-domain Visual Prompts on nuScenes “Seen” Objects
 This visualization focuses on the TTC’s performance with visual prompts in styles different from the training distribution.
-As shown in Figure 18 and Figure 19, the model can effectively detect target objects using visual prompts in various views and styles that diverge from the original training data.
-These results demonstrate the potential of the TTC system to be extended to customized online 3D detection scenarios, where users can provide arbitrary visual prompts to guide the detection of objects of interest.
-The model’s ability to handle prompts across diverse style domains highlights its flexibility and versatility, a key advantage for enabling user-centric, interactive 3D perception systems.
-E.4 Out-domain Visual Prompts on Real-world Examples
+As shown in Figure˜18 and Figure˜19, the model can effectively detect target objects using visual prompts in various views and styles that diverge from the original training data.
+C.4 Out-domain Visual Prompts on Real-world Examples
 This visualization examines the generalization and robustness of TTC detectors in aligning visual prompts with the corresponding target objects in the input video stream.
@@
 This is aimed at demonstrating the TTC system’s capability in reducing online driving risks in such non-standard situations.
-As shown in Figure 20 and Figure 21, our method can effectively localize non-expected animals with higher responses in the regions where animals located, even though it was trained solely on the nuScenes dataset.
-Figure 22, Figure 23, Figure 24, Figure 25 further present examples of detection in a real-world scenario containing both vehicles and animals.
+As shown in Figure˜20 and Figure˜21, our method can effectively localize non-expected animals with higher responses in the regions where animals located, even though it was trained solely on the nuScenes dataset.
+Figure˜22, Figure˜23, Figure˜24, Figure˜25 further present examples of detection in a real-world scenario containing both vehicles and animals.
 Our method can detect all objects with their visual prompts simultaneously.
-This further exemplifies the strong generalization capability of our TTC system, underscoring its potential for effectively handling challenging, edge-case scenarios on roads.
-The ability to accurately detect and localize unexpected objects beyond the training distribution highlights the robustness of the proposed approach, a key requirement for reliable autonomous driving systems.
-Appendix F License of Assets
-The adopted nuScenes dataset [1] is distributed under a CC BY-NC-SA 4.0 license. We implement the model based on mmDet3D codebase [8], which is released under the Apache 2.0 license.
+C.5 License of Assets
+The adopted nuScenes dataset [nuscenes2019] is distributed under a CC BY-NC-SA 4.0 license.
+We implement the model based on mmDet3D codebase [mmdet3d2020], which is released under the Apache 2.0 license.
 We will publicly share our code and models upon acceptance under Apache License 2.0.
@@
 We provide a visual prompt from a separate image as the pre-defined prompt and visualize it at the beginning of the sequence.
-As described, our TTC system can successfully detect the ”Deer” object using this stylized deer prompt sourced from the internet.
+As described, our TTC system can successfully detect the "Deer" object using this stylized deer prompt sourced from the internet.
 This example highlights the TTC’s capability to effectively leverage diverse, user-supplied visual prompts to accurately identify target objects, even when the prompts are not directly from the same scene.
 Figure 25: Visualizations on real world scenarios with out-of-domain visual prompts.
-Another case demonstrates the use of an internet-sourced ”Deer” visual prompt to detect the deer in a different real-world scenario.
+Another case demonstrates the use of an internet-sourced "Deer" visual prompt to detect the deer in a different real-world scenario.
 As shown, our TTC system effectively detects the deer even when it is partially obscured by snow.